### v1

# Closing the Gap on the Sample Complexity of 1-Identification

Abstract
1-identification is a fundamental multi-armed bandit formulation on pure exploration. An agent aims to determine whether there exists a qualified arm whose mean reward is not less than a known threshold $\mu_{0}$, or to output None if it believes such an arm does not exist. The agent needs to guarantee its output is correct with probability at least $1-\delta$, while making expected total pulling times $\mathbb{E}\tau$ as small as possible. We work on 1-identification with two main contributions. (1) We utilize an optimization formulation to derive a new lower bound of $\mathbb{E}\tau$, when there is at least one qualified arm. (2) We design a new algorithm, deriving tight upper bounds whose gap to lower bounds are up to a polynomial of logarithm factor across all problem instance. Our result complements the analysis of $\mathbb{E}\tau$ when there are multiple qualified arms, which is an open problem left by history literature.
Keywords 1-identification, Fixed Confidence, Pure Exploration
1 Introduction
1.1 Model & Problem Formulation
We study the 1-identification problem [^1] This name origins from Katz-Samuels and Jamieson (2020), Li and Cheung (2025). In Degenne and Koolen (2019), it is also called Any Low Problem., a variant of the best arm identification problem. An instance of 1-identification consists of a tuple of four elements: $\nu=([K],\rho=\{\rho_{a}\}_{a\in[K]},\mu_{0},\delta)$. The set $[K]$ denotes the arm set $\{1,2,\cdots,K\}$, and we assume $K\geq 2$. For each $a\in[K]$, $\rho_{a}$ is the reward probability distribution associated with arm $a$. The reward distribution $\rho_{a}$ is not known to the agent for each $a$. By pulling arm $a$ once, the agent receives a random reward distributed as $\rho_{a}$. We denote $r_{a}^{\nu}:=\mathbb{E}_{R\sim\rho_{a}}[R]$ as the mean reward of arm $a$ under reward distribution $\rho_{a}$. We assume the random noise $R_{a}-r_{a}^{\nu}$ to be 1-sub-Gaussian. The real number $\mu_{0}$ is a threshold known to the agent. The agent’s main goal is to predict whether there is an arm $a\in[K]$ with $r_{a}^{\nu}\geq\mu_{0}$, and to identify such an arm in the positive case, while outputting “None” in the negative case. The quantity $\delta\in(0,1)$ is the tolerant parameter on the probability of a wrong prediction.
The agent’s pulling strategy is specified by a sampling rule $\pi=\{\pi_{t}\}_{t=1}^{\infty}$, a stopping time $\tau$ and a recommendation rule $\hat{a}\in[K]\cup\{\textsf{None}\}$. When a pulling strategy is applied on a 1-identification problem instance with reward distribution $\rho$, the strategy generates a history $\{U_{0}\}\cup\{A_{t},X_{t},U_{t}\}_{t=1}^{\tau}$, where $A_{t}\in[K]$ denotes the arm pulled at $t$, $X_{t}\sim\rho_{A_{t}}$ denotes the reward received, and $U_{t}\sim\text{Unif}(0,1)$ is an internal random seed. Arm $A_{t}\in[K]$ is chosen contingent upon the history $H(t-1)=\{U_{0}\}\cup\{(A_{s},X_{s},U_{s})\}^{t-1}_{s=1}$, via the function $\pi_{t}(H(t-1))$. The agent stops pulling at the end of time step $\tau$, where $\tau$ is a stopping time with respect to the filtration $\{\sigma(H(t))\}^{\infty}_{t=1}$. After time step $\tau$, the agent outputs the prediction $\hat{a}\in[K]\cup\{\textsf{None}\}$. We require that $\hat{a}$ is measurable with respect to $\sigma(H(\tau))$, i.e. determined by the information $H(\tau)$. Outputting $\hat{a}\in[K]$ means that the agent concludes with arm $\hat{a}$ satisfying $r_{\hat{a}}^{\nu}\geq\mu_{0}$, while outputting $\hat{a}=\textsf{None}$ means that the agent concludes with no arm having mean reward $\geq\mu_{0}$. We say that $\nu$ is a positive instance if $\max_{a}r_{a}^{\nu}>\mu_{0}$, and say that $\nu$ is a negative instance if $\max_{a}r_{a}^{\nu}<\mu_{0}$.
For $\Delta>0$, we define $\mathcal{S}^{\text{pos}}_{\Delta}=\{\nu:\max_{a}r_{a}^{\nu}-\mu_{0}\geq\Delta\}$, $\mathcal{S}^{\text{neg}}_{\Delta}=\{\nu:\mu_{0}-\max_{a}r_{a}^{\nu}\geq\Delta\}$, and take $\mathcal{S}^{\text{pos}}=\cup_{\Delta>0}\mathcal{S}^{\text{pos}}_{\Delta},\mathcal{S}^{\text{neg}}=\cup_{\Delta>0}\mathcal{S}^{\text{neg}}_{\Delta}$. We define $i^{*}(\nu)$ as the set of correct answers. For $\nu\in\mathcal{S}^{\text{pos}}_{\Delta}$, we have $i^{*}(\nu)=\{a:r_{a}^{\nu}\geq\mu_{0}\}$, while for instance $\nu\in\mathcal{S}^{\text{neg}}_{\Delta}$, we have $i^{*}(\nu)=\{\textsf{None}\}$. We don’t consider instance whose $\max r_{a}^{\nu}=\mu_{0}$ and mainly focus on analyzing PAC algorithms, with the following definitions.
Definition 1.
A pulling strategy is $\delta$-PAC, if for any $\delta\in(0,1),\nu\in\mathcal{S}^{\text{pos}}\cup\mathcal{S}^{\text{neg}}$, it satisfies $\Pr_{\nu}(\tau<+\infty,\hat{a}\in i^{*}(\nu))>1-\delta$.
Appendix A.3 in Li and Cheung (2025) shows that a $\delta$-PAC algorithm can still suffer from $\mathbb{E}_{\nu}\tau=+\infty$ for some $\nu\in\mathcal{S}^{\text{pos}}$. Thus, we further introduce the following definition, which can be considered “Stronger $\delta$-PAC”.
Definition 2.
A pulling strategy is $(\Delta,\delta)$-PAC, if it is $\delta$-PAC, and for any $\Delta,\delta>0$, we have $\sup_{\nu\in\mathcal{S}^{\text{pos}}_{\Delta}\cup\mathcal{S}^{\text{neg}}_{\Delta}}\mathbb{E}_{\nu}\tau<+\infty$.
[…]
1.2 Lit Review on 1-identification
Before illustrating our main contribution, we first review the weakness in the history literature and we start the review by unifying the notations. In the following paper, we consider a reward vector $\{\mu_{a}\}_{a=1}^{K}\subset[0,1]^{K}$ satisfies that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$. Also, we consider $\mu_{0}\in[0,1]$ and define $\Delta_{i,j}=|\mu_{i}-\mu_{j}|$ for $\forall i,j\in[K]\cup\{0\}$. For an instance $\nu$, we assume $r_{a}^{\nu}=\mu_{a}$ without further explanation. In addition, we introduce different complexity terms occurred in the history result.
$\begin{split}H_{1}^{\text{neg}}=&\sum_{a=1}^{K}\frac{2}{\Delta^{2}_{0,a}},H^{\text{low}}_{1}=\sum_{a:\mu_{a}<\mu_{0}}\frac{2}{\Delta_{1,a}^{2}},H=\frac{2}{\Delta^{2}_{0,1}}\\ H_{1}=&\sum_{a=2}^{K}\frac{2}{\Delta^{2}_{1,a}},H_{0}=\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta^{2}_{0,a}}\\ H^{\text{u}}=&\frac{(\frac{K}{m}-1)}{\Delta^{2}},\Delta:=\min_{a:\mu_{a}>\mu_{0}}\Delta_{a,0},m:=\arg\max\{a:\mu_{a}>\mu_{0}\}\\ H(j)=&\frac{1}{j}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{j,a}^{2},\Delta_{a,0}^{2}\}}\end{split}$ (1)
The discussion on the sample complexity focuses on two cases: instance $\nu$ is negative or positive. There has been abundant research achieving the tight approximation of sample complexity on a negative instance, which is similar to the Best Arm Identification. Regarding the lower bound of $\mathbb{E}\tau$, Theorem 5.4 in Li and Cheung (2025) asserts $\mathbb{E}\tau\geq\Omega(H_{1}^{\text{neg}}\log\frac{1}{\delta})$, if we apply any $\delta$-PAC algorithm on an instance $\nu$. Meanwhile, Algorithm HDoC in Kano et al. (2017), APGAI in Jourdan and Réda (2023), SEE in Li and Cheung (2025), are able to guarantee $\mathbb{E}\tau\leq O(H_{1}^{\text{neg}}(\log\frac{1}{\delta}+\log H_{1}^{\text{neg}}))$, ignoring some logarithm factor. In addition, Algorithm Sticky Track and Stop (S-TaS) in Degenne and Koolen (2019) is able to guarantee $\lim\sup\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\leq T^{*}(\nu)$, where $T^{*}(\nu)=H_{1}^{\text{neg}}$ if $\nu$ is negative and the reward follows Gaussian distribution with variance $1$.
Contrast with the negative instance, the complexity of positive instance remains a mystery, as the proposed upper bounds for a positive instance suffer from different looseness. For simplicity, we consider a positive instance $\nu$ and denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and only consider reward following Gaussian distribution with variance $1$. We propose Table 1 to summarize the looseness in the current result, with imprecise formulation because of space limit. The actual upper bound existing in the history literature is not smaller than our presented formulation, and we can still observe the looseness in our presented formulation. Following the Appendix A.2 in Li and Cheung (2025), we put details of simplification on history upper bound in Appendix A.
Table 1: Comparison of non-asymptotic upper bounds on $\mathbb{E}\tau$ for a positive instance $\nu$. “LB” denotes Lower Bound. $\surd_{\text{cond}}$ means the upper or lower bounds are optimal under some specific instances. The definitions of different $H$’s are at the equation (1). We determine whether a sampling complexity upper bound is nearly optimal by comparing with the lower bound. An upper bound is nearly optimal iff the gap is up to a polynomial of $\log K,\{\log\frac{1}{\Delta_{a,0}^{2}}\}_{a=1}^{K}$.
Algorithm Non-asymptotic Bound Optimality
S-TaS, Degenne and Koolen (2019)
Poiani et al. (2025)
$\mathbb{E}\tau\leq O(\frac{\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\Delta_{0,1}^{4}})$ $\times$
HDoC, Kano et al. (2017) $\mathbb{E}\tau\leq O(\frac{\log\frac{K}{\delta}}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}})$ $\times$
lilHDoC, Tsai et al. (2024) $\mathbb{E}\tau\leq O(\frac{\log\frac{K}{\delta}}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}})$ $\times$
APGAI, Jourdan and Réda (2023) $\mathbb{E}\tau\leq O(H_{0}(\log\frac{K}{\delta}))$ $\times$
FWER-FWPD
Katz-Samuels and Jamieson (2020)
$\mathbb{E}\tau\leq O\Big(H^{\text{u}}\log\frac{1}{\delta}\log\Big(H^{\text{u}}\log\frac{1}{\delta}\Big)\Big)$ $\times$
SEE, Li and Cheung (2025) $\mathbb{E}\tau\leq O(\frac{\log\frac{1}{\delta}}{\Delta_{0,1}^{2}}+H(1)\log\frac{K}{\Delta_{0,1}^{2}})$ $\surd_{\text{cond}}$
Parallel-SEE, This work $\mathbb{E}\tau\leq O\Big(\log^{4}K\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)\log\frac{1}{\Delta_{j,0}^{2}}\Big)$ $\surd$
LB, Katz-Samuels and Jamieson (2020) $\mathbb{E}\tau\geq\Omega(\frac{1}{m}H^{\text{low}}_{1}-\frac{1}{\Delta^{2}_{1,m+1}})$ $\times$
LB, Li and Cheung (2025) $\mathbb{E}\tau\geq\Omega(H\log\frac{1}{\delta}+\frac{1}{m}H^{\text{low}}_{1}-\frac{1}{\Delta^{2}_{1,m+1}})$ $\surd_{\text{cond}}$
LB, This work $\mathbb{E}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$ $\surd$
Degenne and Koolen (2019) propose algorithm S-TaS, which provides asymptotic guarantee for both negative and positive instances. S-TaS is able to guarantee $\lim\sup\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\leq\frac{2}{\Delta_{0,1}^{2}}$. Together with their proposed lower bound $\lim\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\geq\frac{2}{\Delta_{0,1}^{2}}$, we know S-TaS is asymptotically optimal. Degenne and Koolen (2019) leaves non-asymptotic analysis of S-TaS as a future direction, which is complemented by Poiani et al. (2025). Poiani et al. (2025) present the non-asymptotic upper bound in an implicit way, but we are able to derive a smaller but explicit formulation. The weakness of Poiani et al. (2025) comes from the $\delta$-independent part $\frac{K^{2}}{\Delta_{0,1}^{4}}$, which is inconsistent with any lower bound.
Algorithm HDoC in Kano et al. (2017) proves an upper bound not smaller than $C^{\text{HDoC}}_{\text{pos}}:=\frac{\log(K/\delta)}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}}$, where $\epsilon=\min\Big\{\min\limits_{a\in[K]}\Delta_{0,a},\min\limits_{a\in[K-1]}\frac{\Delta_{a,a+1}}{2}\Big\}$. The suboptimality of $C^{\text{HDoC}}_{\text{pos}}$ consists of two parts. The first part is $H_{1}\log\log\frac{1}{\delta}$, with comments in Kano et al. (2017): “Therefore, in the case that $\ldots$ $K=\Omega(\frac{\log\frac{1}{\delta}}{\log\log\frac{1}{\delta}})$there still exists a gap between the lower bound in …and the upper bound in …”. The second part is $\frac{K}{\epsilon^{2}}$, which can be infinity if $\mu_{i}=\mu_{j}$ holds. for some $i,j\in[K]$. Based on HDoC, Tsai et al. (2024) propose Algorithm lilHDoC, which incorporate the LIL concentration event (Lemma 3 in Jamieson et al. (2014)) and add an extra warm-up stage. After adding these designs, the proposed algorithm lilHDoC still suffers from the same sub-optimality as HDoC.
Kano et al. (2017) design Algorithm APGAI to address Good Arm Identification, aiming to output all the arms whose reward is above a given threshold $\mu_{0}$. Though it is not the main focus, they also propose an upper bound for 1-identification. The weakness of upper bound $O(H_{0}(\log\frac{K}{\delta}))$ mainly comes from $H_{0}$. Easy to see $H_{0}$ is proportional to $K$, suggesting that the coefficient of $\log\frac{1}{\delta}$ is suboptimal compared to the others. What’s more, if there exists an arm $a$ such that $\mu_{a}=\mu_{0}$, we have $H_{0}=\infty$, suggesting that the upper bound is vacuous.
Katz-Samuels and Jamieson (2020) work on $\epsilon$-good identification and 1-identification simultaneously. They define a new concept Unverifiable Sample Complexity and prove upper bounds for it. Though Unverifiable Sample Complexity is different with Sample Complexity in this paper, we can still adapt their result, as discussed by Appendix A.1 in Li and Cheung (2025). The suboptimality of its upper bound consists of two points. First, its upper bound relies on $\log\frac{1}{\delta}\log\log\frac{1}{\delta}$ instead of $\log\frac{1}{\delta}$. Second, the complexity term $H^{u}$ relies on the minimum gap between a qualified mean reward and the threshold $\mu_{0}$, which can be much smaller than $\Delta_{0,1}$. In the case $m$ is much smaller than $K$ and $\Delta<\Delta_{0,1}$, this upper bound is much larger than their proposed lower bound.
Li and Cheung (2025) design Algorithm SEE, splitting the upper bound of $\mathbb{E}\tau$ into two parts which are $\delta$-dependent and $\delta$-independent. They prove both upper and lower bounds of $\mathbb{E}\tau$ with explicit formulation. They show when $m=1$, i.e. there is only one arm whose reward is greater than $\mu_{0}$, the gap between their proposed upper and lower bounds are up to a polynomial of $\log K,\log\frac{1}{\Delta_{0,1}^{2}}$, suggesting the optimality. However, they also comment $H(1)\log\frac{K}{\Delta_{0,1}^{2}}$, which is $\delta$-independent part, might be improved if $m>1$. They leave the possible improvement as an open direction.
1.3 Main Contribution
In this paper, we develop novel upper and lower bounds for the 1-identification problem. Regarding the lower bound, we conclude
Theorem 1.
For any $\delta$-PAC alg and any reward vector $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, any tolerance level $\delta$ such that $\delta<\min\{\frac{1}{10^{-8}},\frac{1}{64m^{2}}\}$, we can find a permutation $\sigma:[K]\rightarrow[K]$ and define a positive instance $\nu$ with $r_{\sigma(a)}^{\nu}=\mu_{a},a\in[K]$ with unit-variance Gaussian reward, such that $\mathbb{E}_{\text{alg},\nu}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$.
Regarding the upper bound, we merge and adapt the idea in Katz-Samuels and Jamieson (2020) and Li and Cheung (2025) and design a new algorithm Parallel Sequential Exploration Exploitation on Brackets(Algorithm 1, PSEEB). We can prove the following two theorems.
Theorem 2.
Algorithm 1 is $\delta$-PAC.
Theorem 3.
Apply Algorithm 1 to a 1-identification instance $\nu$, we have
$\mathbb{E}_{\nu}\tau\leq\begin{cases}O\left(\log K\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\log^{3}K(\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}})H(j)\right)&\nu\text{ is positive}\\ O\left((\log K)H_{1}^{\text{neg}}\log\frac{H_{1}^{\text{neg}}}{\delta}\right)&\nu\text{ is negative}\end{cases}$
combining Theorem 2 and 3, we know algorithm 1 is $(\Delta,\delta)$-PAC. What’s more, the gap between Theorem 1 and 3 is up to a polynomial of $(\log K,\log\{\frac{1}{\Delta_{0,a}^{2}}\}_{a=1}^{m})$, suggesting that our proposed algorithm 1 is nearly optimal for all 1-identification instances, and $\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)$ is an appropriate complexity measure. Compared to history literature, our proposed Algorithm 1 complement one of the open direction left in Li and Cheung (2025), by deriving a new upper bound for $\mathbb{E}\tau$ when there are multiple qualified arms.
1.4 Literature Review on Other Related Works
If we consider the given threshold $\mu_{0}$ as an arm with deterministic reward, we can address 1-identification by solving a Best Arm Identification(BAI) problem Even-Dar et al. (2002), Gabillon et al. (2012), Jamieson et al. (2013), Kalyanakrishnan et al. (2012), Karnin et al. (2013), Jamieson and Nowak (2014), Chen and Li (2015). However, this is an inefficient solution, suggested by the asymptotic result in Degenne and Koolen (2019), Garivier and Kaufmann (2016) and non-asymptotic result in Li and Cheung (2025), Kano et al. (2017). The reason is 1-identification doesn’t care about an exact comparison between arms with high confidence level, and we should not pull arms as frequently as BAI to distinguish two arms.
Besides 1-identification, there are other variants of pure exploration with multiple answers, as mentioned by Degenne and Koolen (2019). The agent needs to find all the arms Mason et al. (2020) in $\{a:\mu_{a}\geq\mu_{1}-\epsilon\}$ or only one arm Even-Dar et al. (2002), Gabillon et al. (2012), Kalyanakrishnan et al. (2012), Kaufmann and Kalyanakrishnan (2013), Katz-Samuels and Jamieson (2020). As discussed in Katz-Samuels and Jamieson (2020), if the gap $\mu_{1},\mu_{0}$ are known and $\mu_{1}>\mu_{0}$, by taking $\epsilon=\Delta_{0,1}$, the answer set of $\epsilon$-Good Arm Identification and 1-identification are the same. We can adopt same pulling strategy and recommendation rule. But it is unclear about the relationship if there are unknown. What’s more, Locatelli et al. (2016), Mukherjee et al. (2017) work on Thresholding Bandits, which aims to output a set of arms whose reward are above a given $\mu_{0}$, within fixed budget. The optimal set might not be unique under their proposed problem setting.
In this paper, we adopt the idea of parallel algorithm copies, which also exists in the history literature. Chen and Li (2015), Chen et al. (2017), Katz-Samuels and Jamieson (2020) all require a main algorithm to manage different algorithm copies. Katz-Samuels and Jamieson (2020) allocate $2^{\ell}$ time rounds to a copy each time with increasing index $\ell$. Chen and Li (2015), Chen et al. (2017) maintain an overall round index $t\in\mathbb{N}$ and further assign algorithm index to each copy. They execute an algorithm copy only when its algorithm index divide $t$. In our work, we apply round-robin procedure to execute each algorithm copy.
2 Lower Bound of Sample Complexity, for a Positive Instance
In this section, we continue to work on $\{\mu_{a}\}_{a=1}^{K}$ such that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$. We assume each reward distribution follows a unit-variance Gaussian distribution in our lower bound analysis. For an instance $\nu^{\prime}$ and a permutation $\sigma:[K]\rightarrow[K]$, represented as a bijection from $[K]$ to $[K]$, we define the instance $\sigma(\nu^{\prime})$ as $r^{\sigma(\nu^{\prime})}_{\sigma(a)}=r_{a}^{\nu^{\prime}}$.
We first introduce the definition of a symmetric algorithm.
Definition 3.
An algorithm $\widetilde{\text{alg}}$ is said to be symmetric if the following statement holds. For any instance $\nu$ with reward vector $\{\mu_{a}\}_{a=1}^{K}$, any sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$, natural number $T\geq 1$, and two different permutations $\sigma^{\prime},\sigma^{\prime\prime}$, the algorithm $\widetilde{\text{alg}}$ satisfies
$\Pr_{\sigma^{\prime}(\nu),\tilde{\text{alg}}}(\forall t\in[T],A_{t}=\sigma^{\prime}(a_{t}),\tau>T)=\Pr_{\sigma^{\prime\prime}(\nu),\tilde{\text{alg}}}(\forall t\in[T],A_{t}=\sigma^{\prime\prime}(a_{t}),\tau>T).$
Intuitively, under a symmetric algorithm $\widetilde{\text{alg}}$, the reward distribution of the pulled arm $A_{t}$ at time $t$ is invariant under any permutation on the arms’ indexes. In particular, for a symmetric algorithm $\widetilde{\text{alg}}$, we have $\mathbb{E}_{\sigma^{\prime}(\nu),\tilde{\text{alg}}}\tau=\mathbb{E}_{\sigma^{\prime\prime}(\nu),\tilde{\text{alg}}}\tau$ for any two permutations $\sigma^{\prime},\sigma^{\prime\prime}$. A similar definition also appears in Definition 12 of Simchowitz et al. (2017), which further requires the distribution of output $\hat{a}$ is also invariant regarding the distribution. This extra requirement is unnecessary in our analysis. By the following lemma, it suffices to consider only symmetric algorithms for establishing our lower bound.
Lemma 4.
Denote $S$ as the set of all permutations of $[K]$. Given any $\delta$-PAC algorithm alg, we can always find a $\delta$-PAC symmetric algorithm $\widetilde{\text{alg}}$, such that for any instance $\nu^{\prime}$, we have $\max\limits_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau\geq\mathbb{E}_{\tilde{\text{alg}},\nu^{\prime}}\tau$.
Full proof of Lemma 4 is in Appendix B.1. We can shuffle all arms before applying any alg to create a symmetric algorithm $\widetilde{\text{alg}}$. Given Lemma 4, to prove Theorem 1, we suffice to prove the following Theorem.
Theorem 5.
Suppose $\delta<\min\{10^{-8},1/(64m^{2})\}$. For any $\delta$-PAC symmetric algorithm, and for any positive instance $\nu$ with $m\geq 1$ eligible arms, we have
$\mathbb{E}\tau\geq\Omega\left(\min_{1\leq j\leq m}\frac{\log(1/\delta)}{\Delta_{j,0}^{2}}+\frac{1/j}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$
The assumption $\delta<\min\{10^{-8},1/(64m^{2})\}$ is to ensure $\delta\cdot m^{2}$ is upper bounded by a non-trivial absolute constant. The upper bound eliminates the possibility of being $\delta$-PAC by outputting a uniformly chosen arm without any arm pulls.
Sketch Proof of Theorem 5.
In what follows, we present the main steps to prove Theorem 5. The full proof is in Appendix B.2, which consists of the statements and proofs of auxiliary lemmas. The main idea is to establish Lemma 6,
Lemma 6.
Assume $\delta<10^{-8}$, $\frac{1}{8m}\geq\sqrt{\delta}$, we can assert $\mathbb{E}_{\nu}\tau$ is lower bounded by the $\frac{1}{3200}$ times optimal value of the following optimization problem
$\min_{v,\{p_{j}\}_{j=1}^{m}}\quad$ $v$ (2)
s.t. $0\leq p_{j}\leq 1\quad\forall j\in[m]$ (C1)
$v\geq 0$ (C2)
$v\geq\sum_{j=1}^{m}\frac{p_{j}\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}$ (C3)
$\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}$ (C4)
$v\geq\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\quad\forall j\in[m].$ (C5)
If Lemma 6 holds, we first prove the above optimization problem (2) is convex. Then, we turn to construct a feasible solution to the dual problem, and prove the objective value of this feasible solution is lower bounded by $\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$. By the strong duality, we complete the proof of Theorem 5.
Denote event $\mathcal{E}_{j}=\left\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right\}$. To prove Lemma 6, we are going to show taking $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ for $j\in[m]$ is a feasible solution to (2). The overall idea is to apply the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)) to prove the relationship between $\Pr_{\nu}(\mathcal{E}_{j})$ and $\Pr_{\nu^{\prime}}(\mathcal{E}_{j})$ for some alternative instances $\nu^{\prime}$, further the relationship between $3200\mathbb{E}_{\nu}\tau$ and $\{\Pr_{\nu}(\mathcal{E}_{j})\}_{j=1}^{m}$. We prove the feasibility of $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ by going through the constraint (C1), (C2), (C3), (C4) (C5).
For (C1), (C2), they are trivial. As the expected pulling times must be non-negative and the probability must be in the interval $[0,1]$.
For (C3), notice that $\mathds{1}(\tau=+\infty)+\mathds{1}(\tau<+\infty,\hat{a}\in[K]\cup\{\textsf{None}\})=1$, we have $\sum_{j=1}^{m}\mathds{1}(\hat{a}=j)\leq 1$, further $\sum_{j=1}^{m}\mathds{1}(\mathcal{E}_{j})\leq 1$. This observation concludes
$3200\mathbb{E}\tau\geq$ $\sum_{j=1}^{m}\mathbb{E}\tau\mathds{1}(\mathcal{E}_{j})\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\mathcal{E}_{j}).$
For (C4), we start by proving the inequality (3).
$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])\geq 1-5\delta.$ (3)
To prove (3), we consider two cases, $\{a:\mu_{a}=\mu_{0}\}$ is empty or not. The $\delta$-PAC requirement suggests that (3) must hold for both cases.
Then, we are ready to prove (C4) by contradiction. Denote $\bar{\mathcal{E}_{j}}=\{N_{j_{0}}(\tau)<\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}},\hat{a}=j_{0}\}$. If $\sum_{j=1}^{m}\Pr_{\nu}(\mathcal{E}_{j})<\frac{1}{2}$, together with (3), we can find a $j_{0}\in[m]$ such that
$\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)\geq\frac{1}{8m}\geq\sqrt{\delta}.$ (4)
We consider an alternative instance $\nu^{\prime}_{j_{0}}(\epsilon)$ with $r_{a^{\prime}}^{\nu^{\prime}_{j_{0}}}=\begin{cases}\mu_{0}-\epsilon&a^{\prime}=j_{0}\\ r_{a^{\prime}}^{\nu}&else\end{cases}$. The $\delta$-PAC requirement of algorithm conclude
$\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})<\delta$ (5)
Combining (5),(4), we have $\log\frac{\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)}{\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})}\geq\frac{1}{2}\log\frac{1}{\delta}$ holds for all $\epsilon>0$. Meanwhile, Lemma 8, which utilizes the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)), suggests that
$\log\frac{\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)}{\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})}$
$\leq$ $\sqrt{2\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}\log\frac{1}{\sqrt{\delta}}}+\sqrt{2\pi\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}}+\frac{1}{2}\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}$ (6)
Taking $\epsilon\rightarrow 0$, we can show the right hand side of (2) is less than $\frac{1}{2}\log\frac{1}{\delta}$ for $\delta<10^{-8}$, which leads to a contradiction. Lemma 9 merges all the above ideas and conclude $\sum_{j=1}^{m}\Pr_{\nu}(\mathcal{E}_{j})\geq\frac{1}{2}$.
For (C5), we first prove Lemma 11, asserting that
$\mathbb{E}_{\nu}N_{a}(\tau)\geq\frac{1}{4}T_{j}(a)\Pr_{\nu}\left(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right).$ (7)
for each $j\in[m],a\in[K]$, where $T_{j}(a)=\lceil\frac{1}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\Big(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}\Big)}\rceil$. If Lemma 11 holds, we can sum up (7) for each $a\in[K]$ and use the fact that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$, concluding (C5) holds. Then, the remaining work is to prove (7). In the following, we present the sketch idea, which highly relies on the definition of $T_{j}(a)$.
For $n\in\mathbb{N}$, denote
$S_{j,a}^{n}=$ $\{t:N_{j}(t)+N_{a}(t)=n,A_{t+1}\in\{j,a\}\}$
$\tau_{n}^{j}(a)=$ $\min S_{j,a}^{n}$
For any sample path $\{U_{s},A_{s},X_{s}\}_{s=1}^{\tau}$, $S_{j,a}^{n}$ contains at most one element, which is the round that an agent is going to pull arm $a$ or $j$. We suffice to focus on these rounds. In the case that $S_{j,a}^{n}=\emptyset$, we take $\tau_{n}^{j}(a)=+\infty$, which means $\tau_{n}^{j}(a)\geq\tau$. Since $A_{t+1}$ only depends history $H_{t}$, we know $\tau_{n}^{j}(a)$ is a stopping time. Regarding $\tau_{n}^{j}(a)$, we are going to prove two inequalities
$\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)\geq\Pr_{\nu}(\mathcal{E}_{j}),\forall n\leq T_{j}(a).$ (8)
$\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau)\geq\frac{1}{4}\Pr_{\nu}(\tau_{n}^{j}(a)<\tau),\forall n\leq T_{j}(a).$ (9)
If (8) (9) both hold, from the fact that we only pull arm a at $\tau_{n}^{j}(a)+1$, we can conclude
$\mathbb{E}_{\nu}N_{a}(\tau)\geq$ $\sum_{n=0}^{T_{j}(a)}\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<+\infty)$
$\geq$ $\sum_{n=0}^{T_{j}(a)}\frac{1}{4}\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$
$\geq$ $\frac{1}{4}\sum_{n=0}^{T_{j}(a)}\Pr_{\nu}(\mathcal{E}_{j})$
$=$ $\frac{1}{4}T_{j}(a)\Pr_{\nu}(\mathcal{E}_{j}).$
The second step is by (9) and the third step is by (8). Thus, to prove (7), we suffice to show (8) (9) are both correct.
For (8), we notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,0}^{2}}$, further $T_{j}(a)<\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}$. That means for $n\leq T_{j}(a)$, we have $\mathcal{E}_{j}\Rightarrow\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Rightarrow\{\tau_{n}^{j}(a)<\tau\}$, which suggests (8).
For (9), we introduce an alternative instance $\nu_{j,a}$, with $r_{a^{\prime}}^{\nu_{j,a}}=\begin{cases}r_{a}^{\nu}&a^{\prime}=j\\ r_{j}^{\nu}&a^{\prime}=a\\ r_{a^{\prime}}^{\nu}&else\end{cases}$. $\nu_{j,a}$ swaps the position of arm $a$ and $j$ compared to the instance $\nu$. Denote $\tilde{\mathcal{E}}^{aj}(n)=\{\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{j}(n)=\{A_{\tau_{n}^{j}(a)+1}=j,\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{a}(n)=\{A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau\}$. Notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,a}^{2}}$, further $T_{j}(a)\leq\frac{O(1)}{\Delta_{j,a}^{2}(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})})}$, we can use Lemma 7 and 10, which apply the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)), to conclude
$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)$ (10)
where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y}$. To complete this step, we have to put $1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}$ in the denominator of $T_{j}(a)$. (10) suggests that the algorithm is unable to distinguish $\nu$ and $\nu_{ja}$ at the round $t$ such that $N_{a}(t)+N_{j}(t)\leq T_{j}(a)$. For a symmetric algorithm, we have $\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))$, which means
$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)$ (11)
By the Pinsker’s Inequality $kl(x,y)\geq 2(x-y)^{2}$ and $\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j})+\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a})=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj})$, we can derive
$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\geq\frac{1}{4}\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a).$
The above inequality suggests that we we have proved (9), further the correctness of (C5). ∎
3 Upper Bound on the Complexity of 1-identification
In this section, we present our proposed Algorithm PSEEB with rough analysis.
3.1 Algorithm
1Input: Threshold $\mu_{0}$, Error tolerance level $\delta$, Arm number $K$;
2
Randomly generate a permutation of $\sigma:[K]\rightarrow[K]$, take $B_{b}=\{\sigma(i)\}_{i=1}^{\min\{2^{b-1},K\}}$ for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$;
// $\sigma(i)$ denotes the arm index in position $i$
3
4Create algorithm copies $\{\text{alg}_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ by specifying $\text{alg}_{b}$ through calling Algorithm 2 with $B\leftarrow B_{b}$, $\mu_{0}\leftarrow\mu_{0}$, $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$, $C\leftarrow 1.01$, $K\leftarrow K$. Samples across bracket are not shared;
5 for round index $t=1,2,\cdots$ do
6 for Bracket Index $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$ do
7      Execute $\text{alg}_{b}$ for one time round;
8    if $\text{alg}_{b}$ terminates with output $\hat{a}\in[K]\cup\{\textsf{None}\}$ then
9         Terminate the whole algorithm and return $\hat{a}$.
10      end if
11
12   end for
13
14 end for
Algorithm 1 Parallel Sequential Exploration Exploitation on Brackets(PSEEB)
We propose Algorithm 1 to solve the 1-identification problem, which merges and adapts the idea in Katz-Samuels and Jamieson (2020) and Li and Cheung (2025). Katz-Samuels and Jamieson (2020) create each bracket by a unique and independent random shuffle, and a corresponding algorithm copy to pull arms in this bracket. In contrast to them, we specify all the brackets from a single random permutation. Algorithm 1 generates a random permutation of $[K]$ before the start of the pulling process. Based on the random permutation, Algorithm 1 create $\lceil\log_{2}K\rceil+1$ brackets $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$, satisfying $B_{1}\subset B_{2}\subset\cdots\subset B_{\lceil\log_{2}K\rceil+1}=[K]$. For a bracket index $1\leq b\leq\lceil\log_{2}K\rceil$, we have $|B_{b}|=2^{b}$, and $|B_{\lceil\log_{2}K\rceil+1}|=\min\{2^{\lceil\log_{2}K\rceil},K\}$. Our approach is to guarantee $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ follow an increasing order, which is crucial to our Lemma 31. Lemma 31 asserts that with at least some probability, the first several brackets contain at least 1 qualified arm, if there are multiple qualified arms.
Then, Algorithm 1 calls Algorithm 2, defined in Appendix C, as an oracle to set up $\lceil\log_{2}K\rceil+1$ algorithm copies. Algorithm 2 is adapted from the Algorithm 2,3,4,5 in Li and Cheung (2025), named as Sequential Exploration Exploitation(SEE). Algorithm copy $b\in[\lceil\log_{2}K\rceil+1]$ takes bracket $B_{b}$ as an input parameter and only pull arms in $B_{b}$. To fit in the bracketing idea, we modify its budget allocation and stopping rules. Detailed discussion of Algorithm 2 is in Appendix C.
Algorithm 1 manages all the algorithm copies by a round-robin manner. In each round, Algorithm 1 execute one of the algorithm copies with one arm pull and check whether the copy terminates with an output $\hat{a}\in[K]\cup\{\textsf{None}\}$. By the design of Algorithm 2, $\text{alg}_{\lceil\log_{2}K\rceil+1}$ might terminate with taking output $\hat{a}\in[K]\cup\{\textsf{None}\}$. While for bracket $b$ whose size is smaller than $K$, $\text{alg}_{b}$ cannot output None but keep pulling arms, even if the $\text{alg}_{b}$ finds that $\max_{a\in B_{b}}\mu_{a}<\mu_{0}$. The output strategy differs with the algorithm SEE in Li and Cheung (2025).
Once there exists an algorithm copy terminates with an output, Algorithm 1 will adopt the output as the final output and terminate the overall sampling process. We require each algorithm copy to pull arms in a parallel manner without sharing samples. Denote $\tau$ as the total length of algorithm 1, and $\tau(B_{b})$ as the total length of copy $\text{alg}_{b}$. We allow $\tau,\tau(B_{b})$ to take infinities. Since any termination of algorithm copy results in the termination of Algorithm 1, we have $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$, further $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{b})$.
Generally speaking, for a positive instance $\nu$ with $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\mu_{K}$, we wish a bracket $B$ with size smaller than $K$ contains a qualified arm $j\in[m]$. Then, we wish the algorithm copy on this bracket $B$ is able to spend less pulling times to output arm $j$. While for a negative instance $\nu$ with $\mu_{0}>\mu_{1}\geq\mu_{2}\geq\cdots\mu_{K}$, it is possible that all the algorithm copies whose index $b\leq\lceil\log_{2}K\rceil$ never terminates. We only wish the copy $\lceil\log_{2}K\rceil+1$ is able to terminate with finite sampling and output $\hat{a}=\textsf{None}$ when encountering a negative instance.
3.2 Performance Guarantee of Algorithm 1
In the following of this section, we present the sketch proof of Theorem 2 and 3. We use Figure 1 in Appendix C.3 to summarize the overall steps.
Sketch Proof of Theorem 2.
The first step is prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)=0$, whose main idea is very similar to the proof of Theorem 5.2 in Li and Cheung (2025). Same as Li and Cheung (2025), denote $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ as the minimum phase index such that the concentration events of copy $b$ hold, we can prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)\leq\Pr(\max\{\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1},\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}\}=+\infty)=0$.
Given the first step, we are able to conclude $\Pr\Big(\tau<+\infty\Big)=1$, by the fact that $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$. Then, to prove Algorithm 1 is $\delta$-PAC, we suffice to show $\Pr_{\nu}\left(\hat{a}=\textsf{None}\text{ or }\{\hat{a}\in[K],\mu_{\hat{a}}<\mu_{0}\}\right)<\delta$ for a positive instance $\nu$, and $\Pr_{\nu}\left(\hat{a}\in[K]\right)<\delta$ for a negative instance $\nu$.
The proof for both targets is similar. Recall we deliver $\frac{\delta}{\lceil\log_{2}K\rceil+1}$ to each algorithm copy in the Line 1 o Algorithm 1. We are able to prove $\Pr(\text{alg}_{b}\text{ outputs incorrect }\hat{a})<\frac{O(1)\delta}{\lceil\log_{2}K\rceil+1}$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$. Taking the union bound, we complete the proof. ∎
Sketch Proof of Theorem 3.
We split the proof into two cases, which are positive and negative $\nu$.
If the instance $\nu$ is negative, we focus on analyzing $\tau(B_{\lceil\log_{2}K\rceil+1})$, which is very similar to the negative case of Theorem 5.3 in Li and Cheung (2025). By the fact that $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{\lceil\log_{2}K\rceil+1})$, we complete the proof.
If the instance $\nu$ is positive, denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$. Denote $\tilde{T}_{j}(a)=\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, $\tilde{T}_{j}(a;k)=\frac{\log\frac{4K}{\delta_{k}}+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, we are going to show $\mathbb{E}\tau\leq O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\tilde{T}_{j}(a)\right)$ holds for all $j\in[m]$. What’s more, the constant hidden by the Big O notation is independent of $j$. For any given $j\in[m]$, we introduce $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$, which is the minimum bracket index that contains at least one arm not worse than $\mu_{j}$. As we adopt a random permutation to create brackets, $b_{j}$ is a random variable, and we have $\mathbb{E}\tau=\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\mathbb{E}\tau\mathds{1}(b_{j}=b)$. Then, we turn to analyze $O(\log K)\mathbb{E}\tau(B_{b})\mathds{1}(b_{j}=b)$, which is the upper bound of $\mathbb{E}\tau\mathds{1}(b_{j}=b)$.
Denote $\tau^{\text{ee}}(B_{b}),\tau^{\text{et}}(B_{b})$ as the total length of exploration and exploitation periods of algorithm copy $\text{alg}_{b}$ separately. For any arm $a\in B_{b}$, denote $N_{a}^{\text{ee},B_{b}}$ as the pulling times of arm $a$ in the exploration period. Not hard to see $\tau(B_{b})=\tau^{\text{ee}}(B_{b})+\tau^{\text{et}}(B_{b})$. We are going to prove
$\mathbb{E}\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)\Pr(b_{j}=b),$ (12)
$\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)\leq\sum_{a=1}^{K}O(1)\tilde{T}_{j}(a)\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}+O(\log K)|B_{b}|\Pr(b_{j}=b).$ (13)
If (12)(13) are both correct, we can sum up (12)(13) for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$. Lemma 31, 33, 34 prove upper bounds for $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}|B_{b}|\Pr(b_{j}=b)$ and $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}$, which complete the proof of Theorem 3.
To prove (12), we first show the algorithm must terminate no later than the end of phase $L_{b}^{\text{pos}}=\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{O(1)(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$, if $b_{j}=b$. Notice that $\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}=O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$, and Lemma 17 shows $\Pr(\kappa^{\text{ee}}_{b}\geq k)\leq\frac{\pi^{2}\delta_{k-1}}{6}$, $\Pr(\kappa^{\text{et}}_{b}\geq k)\leq\frac{\pi^{2}\delta}{6\alpha_{k-1}}$, we can take expectation on both sides of $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}\mathds{1}(b_{j}=b)$ to complete the proof of (12).
To prove (13), we first show for each $a\in B_{b}$, we have
$N_{a}^{\text{ee},B_{b}}\mathds{1}(b_{j}=b)\leq O(\tilde{T}_{j}(a;L_{b}^{\text{pos}}))(C+1)^{2}\mathds{1}(b_{j}=b).$ (14)
The intuition is as follows. Given the fact that $j\in B_{b}$, once the pulling times of arm $a$ is higher than $O(\tilde{T}_{j}(a;L_{b}^{\text{pos}}))(C+1)^{2}$, copy $\text{alg}_{b}$ can assert either $\text{UCB}_{a}<\mu_{j}$ or $\text{LCB}_{a}>\mu_{0}$, with a constant confidence level. In both cases, $\text{alg}_{b}$ stops pulling arm $a$ in the exploration period. By the facts that $\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)=(|Q|+\sum_{a\in B_{b}}N_{a}^{\text{ee},B_{b}})\mathds{1}(b_{j}=b)$ and elements in $B_{b_{j}}$ are independent of $\kappa^{\text{ee}}_{b_{j}},\kappa^{\text{et}}_{b_{j}}$, we can sum up (14) for $a\in B_{b}$ and derive
$\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)\leq\sum_{a=1}^{K}O(1)\tilde{T}_{j}(a)\Pr(a\in B_{b},b_{j}=b)+O(\log K)|B_{b}|\Pr(b_{j}=b).$
Since $\Pr(a\in B_{b},b_{j}=b)\leq\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}$, and $\Pr(a\in B_{b})=\frac{|B_{b}|}{K}$, we can conclude (13). ∎
Appendix A Detailed Literature Review
A.1 Simplification of History Result
In this section, we only consider Gaussian reward with unit variance. We discuss more details of the upper bounds presented in Table 1. We simplify the result in the history literature with smaller but explicit formulations, for the comparison in Table 1.
A.1.1 Theorem 2 in Poiani et al. (2025)
Theorem 2 in Poiani et al. (2025) presents the Non-asymptotic analysis for algorithm S-TaS. For an instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$, $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and , they define $\neg a=\{\nu^{\prime}:r_{a}^{\nu^{\prime}}<\mu_{0}\}$ for $a\in[m]$, and
$i_{F}(\nu)=\arg\max_{a\in[K]}\sup_{w\in\Delta_{K}}\inf_{\nu^{\prime}\in\neg a}\sum_{a^{\prime}\in K}\frac{w_{a^{\prime}}(r_{a^{\prime}}^{\nu^{\prime}}-r_{a^{\prime}}^{\nu})^{2}}{2}.$
Then, Theorem 2 in Poiani et al. (2025) considers $\epsilon_{\nu}$ such that for all instance $\nu^{\prime}$ such that
$\|\{r_{a}^{\nu^{\prime}}\}_{a=1}^{K}-\{r_{a}^{\nu}\}_{a=1}^{K}\|_{\infty}\leq\epsilon_{\nu}\implies i_{F}(\nu^{\prime})\subset i_{F}(\nu)\cup([K]\setminus i^{*}(\nu)).$
Here $i^{*}(\nu)=[m]$. Let
$T_{\nu}=\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\sqrt{\frac{8D_{K}\log n}{\sqrt{\sqrt{n}+K^{2}}-2K}}\leq\epsilon_{\nu}\right\}\right\}$
where $D_{K}\geq 10$. The upper bound for $\mathbb{E}_{\nu}\tau$ is
$\mathbb{E}_{\nu}\tau\leq 10K^{4}+T_{0}(\delta)+\frac{\pi^{2}}{24}$
where
$T_{0}(\delta)=\inf\left\{t\in\mathbb{N}:\log(Ct^{2}/\delta)\leq(t-\sqrt{t}-1-T_{\nu})\frac{\Delta_{0,1}^{2}}{2}-g(t)\right\}$
and $C\geq e\sum_{t=1}^{+\infty}(\frac{e}{K})^{K}\frac{(\log^{2}(Ct^{2})\log(t))^{K}}{t^{2}}$, $g(t)=\tilde{O}(K^{2}\sqrt{D_{K}t}+\sqrt{D_{K}t^{\frac{3}{2}}})$.
To simplify the above result, we can prove $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\Delta_{0,1}^{4}}$ is a lower bound of $10K^{4}+T_{0}(\delta)+\frac{\pi^{2}}{24}$. From the definition of $T_{\nu}$, we know
$T_{\nu}\geq$ $\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\sqrt{\frac{8D_{K}\log n}{\sqrt{\sqrt{n}+K^{2}}-2K}}\leq\epsilon_{\nu}\right\}\right\}$
$\geq$ $\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\sqrt{\frac{8D_{K}}{\sqrt{\sqrt{n}+K^{2}}-2K}}\leq\epsilon_{\nu}\right\}\right\}$
$=$ $\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\frac{80}{\epsilon_{\nu}^{2}}+2K\leq\sqrt{\sqrt{n}+K^{2}}\right\}\right\}$
$\geq$ $\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\frac{320K}{\epsilon_{\nu}^{2}}+4K^{2}\leq\sqrt{n}+K^{2}\right\}\right\}$
$\geq$ $\max\left\{10K^{4},\frac{90000K^{2}}{\epsilon_{\nu}^{4}}\right\}.$
We can further derive
$T_{0}(\delta)\geq$ $\inf\left\{t\in\mathbb{N}:\log(C/\delta)\leq\left(t-\max\left\{10K^{4},\frac{90000K^{2}}{\epsilon_{\nu}^{4}}\right\}\right)\frac{\Delta_{0,1}^{2}}{2}\right\}$
$=$ $\frac{2\log(C/\delta)}{\Delta_{0,1}^{2}}+\max\left\{10K^{4},\frac{90000K^{2}}{\epsilon_{\nu}^{4}}\right\}.$
Thus, we can conclude $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\epsilon_{\nu}^{4}}$ is smaller than the proposed upper bound in Poiani et al. (2025). The remaining work is to show $\epsilon_{\nu}\leq\Delta_{0,1}$. Denote $a^{\text{sub}}=\arg\max_{a:\mu_{a}<\mu_{1}}\mu_{a}$. We have
  • If $\mu_{a^{\text{sub}}}\geq\mu_{0}$, we assert $\epsilon_{\nu}\leq\Delta_{1,a^{\text{sub}}}<\Delta_{1,0}$. If not, we can set $\nu^{\prime}$ with $r_{a^{\text{sub}}}^{\nu^{\prime}}=\mu_{1}$ while keeping $r_{a}^{\nu^{\prime}}=r_{a}^{\nu^{\prime}}$ for $a\neq a^{\text{sub}}$. With this construction, we have $a^{\text{sub}}\in i_{F}(\nu^{\prime}),a^{\text{sub}}\notin\{1,m+1,m+2,\cdots,K\}$, contradicting to the requirement of $\epsilon_{\nu}$.
  • If $\mu_{a^{\text{sub}}}<\mu_{0}$, we assert $\epsilon_{\nu}\leq\Delta_{1,0}$. If $\epsilon_{\nu}>\Delta_{1,0}$, we can take $\nu^{\prime}$ with $r_{a}^{\nu^{\prime}}=\begin{cases}\mu_{1}-\epsilon_{\nu}&\text{if }\mu_{a}=\mu_{1}\\ \mu_{a}&\text{if }\mu_{a}<\mu_{1}\end{cases}$ for $a\in[K]$. Easy to validate $\|\{r_{a}^{\nu^{\prime}}\}_{a=1}^{K}-\{r_{a}^{\nu}\}_{a=1}^{K}\|_{\infty}\leq\epsilon_{\nu}$ and $i_{F}(\nu^{\prime})=\textsf{None}\notin i_{F}(\nu)\cup([K]\setminus i^{*}(\nu))$
Thus, we have shown $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\Delta_{0,1}^{4}}$ is a lower bound of $10K^{4}+T_{0}(\delta)+\frac{\pi^{2}}{24}$.
A.1.2 Theorem 3 in Kano et al. (2017)
Following Appendix A.2 in Li and Cheung (2025), we can also simplify the Theorem 3 in Kano et al. (2017). By applying the algorithm HDoC, for $\nu\in\mathcal{S}^{\text{pos}}$, we have
$\mathbb{E}_{\nu}\tau\leq n_{1}+\sum_{a=2}^{K}\left(\frac{\log(K\max_{j\in[K]}n_{j})}{2(\Delta_{1,a}-2\epsilon)^{2}}+\delta n_{a}\right)+\frac{K^{2-\frac{\epsilon^{2}}{(\min_{a\in[K]}\Delta_{0,a}-\epsilon)^{2}}}}{2\epsilon^{2}}+\frac{K(5+\log\frac{1}{2\epsilon^{2}})}{4\epsilon^{2}},$
where $n_{a}=\frac{1}{(\Delta_{0,a}-\epsilon)^{2}}\log\left(\frac{4\sqrt{K/\delta}}{(\Delta_{0,a}-\epsilon)^{2}}\log\frac{5\sqrt{K/\delta}}{(\Delta_{0,a}-\epsilon)^{2}}\right)$, $\epsilon=\min\Big\{\min\limits_{a\in[K]}\Delta_{0,a},\min\limits_{a\in[K-1]}\frac{\Delta_{a,a+1}}{2}\Big\}$. From direct calculation, we can observe
$n_{1}\geq\frac{\log\frac{K}{\delta}}{2\Delta_{0,1}^{2}},\quad\frac{K(5+\log\frac{1}{2\epsilon^{2}})}{4\epsilon^{2}}\geq\frac{K}{4\epsilon^{2}}$
$\log(K\max_{j\in[K]}n_{j})\geq\log(\frac{K}{2}\log\frac{1}{\delta})\geq\log\log\frac{1}{\delta}.$
From the above inequality, we can conclude
$\sum_{a=2}^{K}\frac{\log(K\max_{j\in[K]}n_{j})}{2(\Delta_{1,a}-2\epsilon)^{2}}$
$\geq$ $\sum_{a=2}^{K}\frac{\log\log\frac{1}{\delta}}{2(\Delta_{1,a}-2\epsilon)^{2}}$
$\geq$ $\sum_{a=2}^{K}\frac{\log\log\frac{1}{\delta}}{2\Delta_{1,a}^{2}}$
$\geq$ $\frac{H_{1}}{2}\log\log\frac{1}{\delta}.$
Thus, we can conclude
$n_{1}+\sum_{a=2}^{K}\left(\frac{\log(K\max_{j\in[K]}n_{j})}{2(\Delta_{1,a}-2\epsilon)^{2}}+\delta n_{a}\right)+\frac{K^{2-\frac{\epsilon^{2}}{(\min_{a\in[K]}\Delta_{0,a}-\epsilon)^{2}}}}{2\epsilon^{2}}+\frac{K(5+\log\frac{1}{2\epsilon^{2}})}{4\epsilon^{2}}$
$\geq$ $\frac{\log\frac{K}{\delta}}{2\Delta_{0,1}^{2}}+\frac{H_{1}}{2}\log\log\frac{1}{\delta}+\frac{K}{4\epsilon^{2}}.$
A.1.3 Thereom 2 in Jourdan and Réda (2023)
Jourdan and Réda (2023) do not explicitly provide upper bound for $\mathbb{E}\tau$. Therorem 2 in Jourdan and Réda (2023) suggests that for a positive instance $\nu$.
$\mathbb{E}_{\nu}\tau$ $\leq C^{\text{pos}}(\delta)+\frac{K\pi^{2}}{6}+1$
where
$C^{\text{pos}}(\delta)=$ $\sup\left\{t:t\leq\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)(\sqrt{c(t,\delta)}+\sqrt{3\log t})^{2}+D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})\right\}$
$2c(t,\delta)=$ $\bar{W}_{-1}\left(2\log\frac{K}{\delta}+4\log\log(e^{4}t)+\frac{1}{2}\right)$
$\bar{W}_{-1}(x)=$ $-W_{-1}(-e^{-x})\approx x+\log x.$
$W_{-1}$ is the negative branch of the Lambert W function, and $D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})$ is further defined in the Appendix F in Jourdan and Réda (2023), as
$D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})=$ $S_{\mu}+1-\frac{6\log S_{\mu}}{\Delta_{0,1}^{2}}$
$S_{\mu}=$ $h_{1}\left(12\sum_{a=1}^{K}\frac{1}{\Delta_{0,a}^{2}},n_{0}K+2m\right)$
$h_{1}(z,y)=$ $z\bar{W}_{-1}(\frac{y}{z}+\log z)$
$\bar{W}_{-1}(x)=$ $-W_{-1}(-e^{-x})$
From the definition, we know
$C^{\text{pos}}(\delta)\geq$ $\sup\left\{t:t\leq\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)c(t,\delta)\right\}$
$\geq$ $\sup\left\{t:t\leq\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\left(\log\frac{K}{\delta}+2\log\log(e^{4}t)+\frac{1}{4}\right)\right\}$
$\geq$ $\sup\left\{t:t\leq\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\log\frac{K}{\delta}+2\log 4\right\}$
$\geq$ $\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\log\frac{K}{\delta}.$
Thus, we can conclude $\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\log\frac{K}{\delta}$ is a lower bound of $C^{\text{pos}}(\delta)+\frac{K\pi^{2}}{6}+1$.
A.1.4 Theorem 8 in Katz-Samuels and Jamieson (2020)
As discussed in the Appendix A.1 in Li and Cheung (2025), proposed Algorithm FWER-FWPD only work on the positive instance with a different definition of $\mathbb{E}\tau$. But we can adapt their result. Apply Algorithm FWER-FWPD on a positive instance $\nu$, Theorem 8 in Katz-Samuels and Jamieson (2020) suggests that
$\mathbb{E}\tau\leq O\left(\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)\log\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)\right),$
where $\Delta=\min_{a:\mu_{a}>\mu_{0}}\Delta_{a,0}$.
Not hard to see
$\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)\log\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)$
$\geq$ $\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}\right)\log\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}\right).$
A.2 Notation Description
Table LABEL:tab:notation presents part of the notation description used in the following appendix.
Table 2: Notation Summary
Symbol Type Meaning
$K$ Integer Number of arms.
$a,i,j$ Integer in $[K]$ Arm Index
$\delta$ Real Tolerance level
$\hat{a}$ NA Output of Algorithm 1, 2, takes value in $K]\cup\{\textsf{None}\}$
$\hat{a}_{k}$ NA Rough guess of an exploration period in Algorithm 2, takes value in $K]\cup\{\textsf{None},\textsf{Not Complete}\}$
$k$ Integer Phase index in Algorithm 2
$\nu,\nu^{\prime},\nu^{\prime\prime},\tilde{\nu}$ NA An 1-identification instance
$\rho_{a}$ NA Reward distribution of arm $a$
$\sigma(\nu)$ NA Switch the arm positions in instance $\nu$ with a permutation $\sigma$
$\mu_{a}$ Real Mean value of arm $a$.
$m$ Integer Maximum arm index whose mean value is above $\mu_{0}$, we usually assume $\mu_{1}\geq\cdots\geq\mu_{m}\geq\mu_{0}>\mu_{m+1}\geq\cdots\geq\mu_{K}$
$r_{a}^{\nu}$ Real Mean value of arm $a$ instance $\nu$.
$\Delta_{i,j}$ Real Mean gap between arm index $i$ and $j$.
$A_{t}$ Integer in $[K]$ Action in round $t$
$B$ Subset of $[K]$ A bracket
$b,\tilde{b}$ Integer Bracket index in $\lceil\log_{2}K\rceil+1]$
$\tau,\tau(B)$ Random Variable $\tau$ is the termination round of Algorithm 1, $\tau(B)$ is the termination round of an Algorithm 2
$\tau_{k}^{\text{ee}}(B),\tau_{k}^{\text{et}}(B)$ Random Variable The total length of the exploration / exploitation periods at to the end of phase $k$, of Algorithm 2 with bracket $B$ as input
$\tau,\tau_{n},\tilde{\tau},\tau^{\prime},\tau_{n}^{j}(a)$ Random Variable Stopping times for analysis
$\mathcal{E}$ Random Event Random Event
$Z_{s},Z_{t}$ Random Variable Realized KL divergence, usually follows $N(\frac{\Delta^{2}}{2},\Delta^{2})$ as we only consider Gaussian distribution with variance $1$ as reward distribution in the proof lower bound
$D,D_{i}$ Random Variable $D:=\max_{1\leq t\leq n}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}}{2}\right)$. See the proof of Lemma 7
$N_{a}(t)$ Random Variable Pulling times of arm $a$ up to round $t$
$T_{j}(a)$ Real Value A pulling times threshold dependent on arm $a$ and $j$
$U(t,\delta)$ Real Value Half length of confidence interval with pulling times $t$
$\mathcal{H}^{ee},\mathcal{H}^{et},Q$ Random Set Collected history in Algorithm 2
$N_{a}^{\text{ee}}(\mathcal{H}^{ee}),N_{a}^{\text{et}}(\mathcal{H}^{et})$ Random Variable Pulling times of arm $a$ in history $\mathcal{H}^{ee},\mathcal{H}^{et}$
$N_{a}^{\text{ee}}(t)$ Random Variable Pulling times of arm $a$ at round $t$ of an exploration period
$N_{a}^{\text{et}}(t)$ Random Variable Pulling times of arm $a$ at round $t$ of an exploitation period
$\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee}),\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})$ Random Variable Empirical mean value of arm $a$ in history $\mathcal{H}^{ee},\mathcal{H}^{et}$
$\text{UCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta)$ Random Variable Upper Confidence Bound calculated from history $\mathcal{H}^{ee}$
$\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta)$ Random Variable Lower Confidence Bound calculated from history $\mathcal{H}^{ee}$
$\text{UCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)$ Random Variable Upper Confidence Bound calculated from history $\mathcal{H}^{et}$
$\text{LCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)$, Random Variable Lower Confidence Bound calculated from history $\mathcal{H}^{et}$
$C$ Real Value Fixed Constant In Algorithm 2, we take $C$ to any real number greater than 1
$t^{\text{ee}},t^{\text{et}}$ Integer Counter of the total pulling times in exploration and exploitation periods, used in Algorithm 2
$\delta_{k},\beta_{k},\alpha_{k}$ Real Value Hyperparameters of Algorithm 2 to control the pulling process
$X_{a,s}^{\text{ee}}$ Random Variable Realized reward by pulling arm $a$ at the $s$-th times in the exploration period
$X_{a,s}^{\text{et}}$ Random Variable Realized reward by pulling arm $a$ at the $s$-th times in the exploitation period
$\kappa^{\text{ee}}_{b},\kappa^{\text{ee}}_{B}$ Random Variable Minimum phase index such that the concentration event holds in the exploration period of algorithm copy $b$. If we only discuss a given bracket $B$, we do not emphasize its bracket index and may use $\kappa^{\text{ee}}_{B}$.
$\kappa^{\text{et}}_{b},\kappa^{\text{et}}_{B}$ Random Variable Minimum phase index such that the concentration event holds in the exploitation period of algorithm copy $b$. If we only discuss a given bracket $B$, we do not emphasize its bracket index and may use $\kappa^{\text{et}}_{B}$.
$\hat{\mu}^{ee}_{a,b}(t)$ Random Variable The empirical mean value of arm $a$ of the first $t$ collected reward in the exploration period of algorithm copy $b$
$\hat{\mu}^{et}_{a,b}(t)$ Random Variable The empirical mean value of arm $a$ of the first $t$ collected reward in the exploitation period of algorithm copy $b$
$\hat{\mu}_{a,t}$ Random Variable The empirical mean value of arm $a$ of the first $t$ collected reward. We use this notation if a specific bracket and a specific exploration/exploitation period is given, we may turn to use $\hat{\mu}_{a,t}$ instead of $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$.
Appendix B Proof of Lower Bound
B.1 Proof of Preparation
The first step is to prove Lemma 4, showing that we suffice to only consider symmetric algorithm
Proof of Lemma 4.
Given the alg, we define $\widetilde{\text{alg}}$ as the following. Shuffle all the arms randomly before the pulling process, and then apply the algorithm alg. We firstly show $\widetilde{\text{alg}}$ is indeed symmetric. For any pulling sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$ and any permutation $\sigma$, we have
$\Pr_{\widetilde{\text{alg}},\sigma(\nu^{\prime})}(\forall t\in[T],A_{t}=a,\tau>t)=\frac{\sum_{\tilde{\sigma}\in S}\Pr_{\text{alg},\tilde{\sigma}(\nu^{\prime})}(\forall t\in[T],A_{t}=a,\tau>t)}{K!},$
which means $\tilde{\text{alg}}$ fits in the definition. Then, we have
$\max_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau$
$\geq$ $\frac{\sum_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau}{K!}$
$\geq$ $\frac{\sum_{\sigma\in S}\sum_{t=1}^{+\infty}\Pr_{\text{alg},\sigma(\nu^{\prime})}(\tau\geq t)}{K!}$
$=$ $\mathbb{E}_{\widetilde{\text{alg}},\nu^{\prime}}\tau$
The last thing is to show $\widetilde{\text{alg}}$ is also $\delta$-PAC. Since alg is $\delta$-PAC, we know for any permutation $\sigma$, we have
$\Pr_{\text{alg},\sigma(\nu)}(\tau<+\infty,\hat{a}\in i^{*}(\sigma(\nu)))>1-\delta.$
That means
$\Pr_{\widetilde{\text{alg}},\sigma(\nu)}(\tau<+\infty,\hat{a}\in i^{*}(\sigma(\nu)))$
$=$ $\frac{\sum_{\sigma\in S}\Pr_{\text{alg},\sigma(\nu)}(\tau<+\infty,\hat{a}\in i^{*}(\sigma(\nu)))}{K!}$
$>$ $\frac{\sum_{\sigma\in S}1-\delta}{K!}$
$=$ $1-\delta.$
We can conclude $\widetilde{\text{alg}}$ is also $\delta$-PAC. ∎
The next step, which is also the main trick, is to the prove Lemma 7.
Lemma 7.
Given a pair $i,j\in[K],i\neq j$ and an arbitrary instance $\nu^{\prime}$, denote $\Delta=|r^{\nu^{\prime}}_{i}-r^{\nu^{\prime}}_{j}|$. Define
$\tau_{n}=\min\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}.$
Apply a symmetric algorithm, we have
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}+$
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\frac{1}{2}n\Delta^{2}\Pr_{\nu^{\prime}}(\tau_{n}<\tau).$
An observation is given the realized reward and a symmetric algorithm, the set $\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}=:S_{ij}(n)$ contains at most one element. If $|S_{ij}(n)|=1$, $\tau_{n}$ will be the unique element, and we have $\tau_{n}<\tau$ holds in this case. If $|S_{ij}(n)|=0$, we have $\tau_{n}=\infty$, meaning that $\tau_{n}\geq\tau$. Another observation is
$\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}+\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}$
is the kl-divergence between two Bernoulli distribution with probability $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)$, $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)$, and not hard to see $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)+\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)=1$.
Proof of Lemma 7.
Denote $\nu^{\prime\prime}$ as the instance which switch the position of arm $i$ and arm $j$ in $\nu^{\prime}$, i.e. $r_{i}^{\nu^{\prime}}=r_{j}^{\nu^{\prime\prime}},r_{j}^{\nu^{\prime}}=r_{i}^{\nu^{\prime\prime}}$ and $r_{a}^{\nu^{\prime}}=r_{a}^{\nu^{\prime\prime}},a\neq i,j$.
We first derive
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$ (15)
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$ (16)
where $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta^{2}}{2},\Delta^{2})$. Since these two inequalities are symmetric, we suffice to prove the first one. Take $\mathcal{E}=\{\tau_{n}<\tau,A_{\tau_{n}+1}=i\}$, from the Lemma 18 in Kaufmann et al. (2016), we have
$\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})+$
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\neg\mathcal{E}\right]\Pr_{\nu^{\prime}}(\neg\mathcal{E})$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})$
$\geq$ $\exp\left(\mathbb{E}_{\nu^{\prime}}\left[-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}|\mathcal{E}\right]\right)\Pr_{\nu^{\prime}}(\mathcal{E}),$
where $\hat{KL}_{i,s}:=\log\frac{\exp(-\frac{(X_{i,s}-r_{i}^{\nu^{\prime}})^{2}}{2})}{\exp(-\frac{(X_{i,s}-r_{i}^{\nu^{\prime\prime}})^{2}}{2})}=\frac{(2X_{i,s}-r_{i}^{\nu^{\prime}}-r_{j}^{\nu^{\prime}})(r_{i}^{\nu^{\prime}}-r_{j}^{\nu^{\prime}})}{2}$, $X_{i,s}\sim N(r_{i}^{\nu^{\prime}},1)$, $\hat{KL}_{j,s}:=\log\frac{\exp(-\frac{(X_{j,s}-r_{j}^{\nu^{\prime}})^{2}}{2})}{\exp(-\frac{(X_{j,s}-r_{j}^{\nu^{\prime\prime}})^{2}}{2})}=\frac{(2X_{j,s}-r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})}{2}$, $X_{j,s}\sim N(r_{j}^{\nu^{\prime}},1)$. The last line is by the Jensen’s Inequality. Following the last line, we have
$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}|\mathcal{E}\right]$
$=$ $\frac{\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)\right]}{\Pr_{\nu^{\prime}}(\mathcal{E})},$
and we can conclude $\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\geq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)\right]$. From the above calculation, we know $\hat{KL}_{i,s},\hat{KL}_{j,s}$ both follow $N(\frac{(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})^{2}}{2},(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}}))$, and we can rewrite the conclusion as
$\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right]$
where $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta^{2}}{2},\Delta^{2})$. The above inequality suggests that we have proved (15) and (16).
By the fact that the algorithm is symmetric, we can conclude $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)$ and $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)$. Sum up inequality (15) and (16), we get
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}+$
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}$
$=$ $\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}+$
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right).$
The remaining work is to prove
$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\frac{1}{2}n\Delta^{2}\Pr_{\nu^{\prime}}(\tau_{n}<\tau).$
Denote $D:=\max_{1\leq t\leq n}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}}{2}\right)$, we have
$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\max_{1\leq t\leq n}\sum_{s=1}^{n}Z_{s}\right)$
$=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\max_{1\leq t\leq n}\sum_{s=1}^{n}Z_{s}-\frac{\Delta^{2}}{2}+\frac{\Delta^{2}}{2}\right)$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\max_{1\leq t\leq n}\sum_{s=1}^{n}Z_{s}-\frac{\Delta^{2}}{2}\right)+\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\frac{n\Delta^{2}}{2}$
$=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)D+\frac{n\Delta^{2}}{2}\Pr_{\nu^{\prime}}(\tau_{n}<\tau).$
The remaining work is to prove $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)D\leq\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)$. To simplify the notation, we take event $\mathcal{E}=\{\tau_{n}<\tau\}$ here and we can conduct the following calculation.
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D\right]$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D\right]$
$=$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathcal{E},D>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\Pr_{\nu^{\prime}}\Big(D>s\Big)\right\}ds$
From Lemma 32, we can conclude $\Pr_{\nu^{\prime}}\big(D\geq s\big)\leq\exp(-\frac{s^{2}}{2n\Delta^{2}})$. Plug in the inequality, we get
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D\right]\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\exp(-\frac{s^{2}}{2n\Delta^{2}})\right\}ds$
$=$ $\int_{0}^{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}\Pr_{\nu^{\prime}}(\mathcal{E})ds+\int_{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2n\Delta^{2}})ds$
$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\frac{\sqrt{2\pi}\sqrt{n\Delta^{2}}}{\sqrt{2\pi}}\int_{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2})ds$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n\Delta^{2}}\exp(-\frac{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}{2})$
$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\mathcal{E}).$
The second last step is by the equality $\int_{t}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{s^{2}}{2}}ds\leq e^{-\frac{t^{2}}{2}}$ for $t>0$. The reason is
$\int_{t}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{s^{2}}{2}}ds=\int_{0}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{(s+t)^{2}}{2}}ds\leq e^{-\frac{t^{2}}{2}}\int_{0}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{s^{2}}{2}}ds\leq e^{-\frac{t^{2}}{2}}.$
We complete the proof. ∎
Besides Lemma 7, we further introduce a Lemma which is similar, but focuses on the single side.
Lemma 8.
Consider two instances $\nu^{\prime}$, $\nu^{\prime\prime}$, which only differ in arm $a_{1},a_{2},\cdots,a_{M}$, i.e. $r^{\nu^{\prime}}_{a}\neq r^{\nu^{\prime\prime}}_{a}$ holds for all $a\in\{a_{1},a_{2},\cdots,a_{M}\}$ and $r^{\nu^{\prime}}_{a}=r^{\nu^{\prime\prime}}_{a}$ holds for all $a\notin\{a_{1},a_{2},\cdots,a_{M}\}$. Denote $\Delta_{i}=|r_{a_{i}}^{\nu^{\prime}}-r_{a_{i}}^{\nu^{\prime\prime}}|$, and $\tilde{\tau}$ is an arbitrary stopping time. If event $\mathcal{E}\in\mathcal{F}_{\tilde{\tau}}$ fulfills $\mathcal{E}\subset\{N_{a_{i}}(\tilde{\tau})\leq n_{i},\forall i\in[M]\}$ with some $\{n_{i}\}_{i=1}^{M}\in\mathbb{N}^{M}$, we can conclude
$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\sum_{i=1}^{M}\sqrt{2n_{i}\Delta_{i}^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}+\sqrt{2\pi n_{i}\Delta_{i}^{2}}+\frac{1}{2}n_{i}\Delta_{i}^{2}$
Proof.
From the Lemma 18 in Kaufmann et al. (2016), we have
$\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})+$
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)|\neg\mathcal{E}\right]\Pr_{\nu^{\prime}}(\neg\mathcal{E})$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})$
$\geq$ $\exp\left(\mathbb{E}_{\nu^{\prime}}\left[-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}|\mathcal{E}\right]\right)\Pr_{\nu^{\prime}}(\mathcal{E}),$
where $\hat{KL}_{i,s}:=\log\frac{\exp(-\frac{(X_{i,s}-r_{a_{i}}^{\nu^{\prime}})^{2}}{2})}{\exp(-\frac{(X_{i,s}-r_{a_{i}}^{\nu^{\prime\prime}})^{2}}{2})}=\frac{(2X_{i,s}-r_{a_{i}}^{\nu^{\prime}}-r_{a_{i}}^{\nu^{\prime\prime}})(r_{a_{i}}^{\nu^{\prime}}-r_{a_{i}}^{\nu^{\prime\prime}})}{2}$, $X_{i,s}\sim N(r_{a_{i}}^{\nu^{\prime}},1)$. Following the last line, we have
$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}|\mathcal{E}\right]$
$=$ $\frac{\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right]}{\Pr_{\nu^{\prime}}(\mathcal{E})}.$
and we can conclude $\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\sum_{i=1}^{M}\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right]$. We turn to find an upper bound for each $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right]$. Since $\mathcal{E}\Rightarrow N_{a_{i}}(\tilde{\tau})<n_{i}$, we have
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right]$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\max_{1\leq t\leq n_{i}}\sum_{s=1}^{t}\hat{KL}_{i,s}-\frac{\Delta_{i}^{2}}{2}\right)\right]+\frac{n_{i}\Delta_{i}^{2}}{2}\Pr_{\nu^{\prime}}(\mathcal{E})$
Denote $D_{i}:=\max_{1\leq t\leq n_{i}}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}_{i}}{2}\right)$, $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta_{i}^{2}}{2},\Delta_{i}^{2})$, we have
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D_{i}\right]$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D_{i}\right]$
$=$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathds{1}(\mathcal{E})\mathds{1}(D_{i}\geq 0)D_{i}>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathcal{E},D_{i}>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\Pr_{\nu^{\prime}}\Big(D_{i}>s\Big)\right\}ds$
From Lemma 32, we can conclude $\Pr_{\nu^{\prime}}\big(D_{i}\geq s\big)\leq\exp(-\frac{s^{2}}{2n_{i}\Delta^{2}_{i}})$. Plug in the inequality, we get
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D_{i}\right]\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\exp(-\frac{s^{2}}{2n_{i}\Delta^{2}_{i}})\right\}ds$
$=$ $\int_{0}^{\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}\Pr_{\nu^{\prime}}(\mathcal{E})ds+\int_{\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2n_{i}\Delta^{2}})ds$
$=$ $\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\frac{\sqrt{2\pi}\sqrt{n_{i}\Delta^{2}_{i}}}{\sqrt{2\pi}}\int_{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2})ds$
$\leq$ $\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n_{i}\Delta^{2}_{i}}\exp(-\frac{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}{2})$
$=$ $\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n_{i}\Delta^{2}_{i}}\Pr_{\nu^{\prime}}(\mathcal{E}).$
In summary, we have proved
$\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}$
$\leq$ $\sum_{i=1}^{M}\sqrt{2n_{i}\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n_{i}\Delta^{2}}\Pr_{\nu^{\prime}}(\mathcal{E})+\frac{n_{i}\Delta_{i}^{2}}{2}\Pr_{\nu^{\prime}}(\mathcal{E}),$
which is equivalent to the conclusion we need to prove. ∎
B.2 Proof of Theorem 5
Following the sketch proof in section 2, we prove the lemmas one by one. The first step is to build a “constant probability lower bound”.
Lemma 9.
Apply a $\delta$-PAC and symmetric algorithm to a positive instance $\nu$, If $\max_{i,j\in[K]}\Delta_{i,j}<1$ and $\delta<10^{-8}$, $\frac{1}{8m}\geq\sqrt{\delta}$, all holds, we have
$\sum_{j=1}^{m}\Pr_{\nu}\Big(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)\geq\frac{1}{2}$
Proof of Lemma 9.
Recall the notation $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, and the $\delta$-PAC assumption requires
$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m]\cup\{a:\mu_{a}=\mu_{0}\})>1-\delta$
[…]
We prove the claim by contradiction, if $Pr_{\nu}(\tau<+\infty,\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 4\delta$ holds, we can find a $T_{0}$, such that
$\Pr_{\nu}(\tau<T_{0},\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 2\delta.$ (17)
We define $\nu^{\prime}(\epsilon)$ with $r_{a}^{\nu^{\prime}(\epsilon)}=\begin{cases}\mu_{0}-\epsilon&\mu_{a}=\mu_{0}\\ r_{a}^{\nu}&\text{else}\end{cases}$, where $\epsilon$ is an arbitrary positive number. We know $i^{*}(\nu^{\prime}(\epsilon))=[m]$ and we have
$\Pr_{\nu^{\prime}(\epsilon)}(\tau<+\infty,\hat{a}\in[m])>1-\delta$
$\Rightarrow$ $\Pr_{\nu^{\prime}(\epsilon)}(\tau=+\infty\text{ or }\hat{a}\notin[m])<\delta$
[…]
$\log\frac{\Pr_{\nu}(\mathcal{E})}{\Pr_{\nu^{\prime}(\epsilon)}(\mathcal{E})}\leq$ $\sum_{a:\mu_{a}=\mu_{0}}\sqrt{2T_{0}\epsilon^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}+\sqrt{2\pi T_{0}\epsilon^{2}}+\frac{1}{2}T_{0}\epsilon^{2}$
$\log\frac{\Pr_{\nu}(\mathcal{E})}{\Pr_{\nu^{\prime}(\epsilon)}(\mathcal{E})}\geq$ $\log\frac{2\delta}{\delta}=\log 2$
[…]
$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m]\cup\{a:\mu_{a}=\mu_{0}\})>1-\delta$
[…]
$\Rightarrow$ $\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])>1-\delta-4\delta$
$\Rightarrow$ $\Pr_{\nu}(\tau=+\infty\text{ or }\hat{a}\notin[m])<5\delta$
$\Rightarrow$ $\Pr_{\nu}(\hat{a}\notin[m])<5\delta$
Now, we are ready to prove the Lemma. We prove the conclusion by contradiction. If the Lemma doesn’t hold, we have
$\Pr_{\nu}\Big(\exists j\in[m],\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)<\frac{1}{2}$
$\Leftrightarrow$ $\Pr_{\nu}\left(\cap_{j=1}^{m}\Big(\{\hat{a}\neq j\}\cup\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\right)\geq\frac{1}{2}$
$\Rightarrow$ $\Pr_{\nu}\left(\cap_{j=1}^{m}\{\hat{a}\neq j\}\right)+\Pr_{\nu}\left(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\right)\geq\frac{1}{2}$
$\Rightarrow$ $5\delta+\Pr_{\nu}\left(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\right)\geq\frac{1}{2}$
$\Rightarrow$ $\Pr_{\nu}\left(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\right)\geq\frac{1}{4}$
$\Rightarrow$ $\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}\in[m]\}\right)+\Pr_{\nu}\left(\{\hat{a}\notin[m]\}\right)\geq\frac{1}{4}$
$\Rightarrow$ $\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}\in[m]\}\right)+5\delta\geq\frac{1}{4}$
$\Rightarrow$ $\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}\in[m]\}\right)\geq\frac{1}{8}$
$\Rightarrow$ $\sum_{j^{\prime}=1}^{m}\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}=j^{\prime}\}\right)\geq\frac{1}{8}$
From the last step, we know there exists $j_{0}$, such that
$\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}=j_{0}\}\right)\geq\frac{1}{8m}\geq\sqrt{\delta}.$
We can consider instance $\nu_{j_{0}}^{(0)}$ whose reward of arm $j_{0}$ is $\mu_{0}-\epsilon$ for some $\epsilon>0$, others remain the same as $\nu$. By the Lemma 8, we have
$\log\frac{\sqrt{\delta}}{\delta}$
$\leq$ $\log\frac{\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}=j_{0}\}\right)}{\Pr_{\nu_{j_{0}}^{(0)}}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}=j_{0}\}\right)}$
$\leq$ $\sqrt{2\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}+\sqrt{2\pi\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}}+\frac{1}{2}\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}$
$\leq$ $\sqrt{2\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}\log\frac{1}{\sqrt{\delta}}}+\sqrt{2\pi\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}}+\frac{1}{2}\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}.$
Taking $\epsilon\rightarrow 0$, the last line suggests that
$\frac{1}{2}\log\frac{1}{\delta}\leq\sqrt{\frac{\log\frac{1}{\delta}}{100}\log\frac{1}{\delta}}+\sqrt{2\pi\frac{\log\frac{1}{\delta}}{100}}+\frac{1}{2}\frac{\log\frac{1}{\delta}}{100},$
which leads to a contradiction, as $\delta<10^{-8}$. ∎
The next step is to prove a lemma analyzing the pulling behavior.
Lemma 10.
Assume $\delta<\frac{1}{e}$. For any $j\in[m]$, $a\in[K],a\neq j$, define
$\tau_{n}^{j}(a)=$ $\min\{t:N_{j}(t)+N_{a}(t)=n,A_{t+1}\in\{j,a\}\}$
$T_{j}(a)=$ $\lceil\frac{1}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}\right)}\rceil.$
Then, for $n=0,1,2,\cdots,T_{j}(a)-1$, we have
$\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=j)\log\frac{\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=j)}{\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=a)}+$
$\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=a)\log\frac{\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=a)}{\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=j)}$
$<$ $0.4\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$
Proof of Lemma 10.
The main idea is to utilize Lemma 7. From the definition, we know $T_{j}(a)\geq 1$. For $n=1,2,\cdots,T_{j}(a)-1$, we have
$\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu}(N_{a}(\tau)+N_{j}(\tau)>n)}}$
$\leq$ $\sqrt{2\Big(\frac{2}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\log\frac{e}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}}\Big)\Delta_{a,j}^{2}\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}}$
$\leq$ $\sqrt{\frac{1}{50}}.$
The third step is by the fact that $\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\}\subset\{N_{a}(\tau)+N_{j}(\tau)>n\}$ holds for all $n=0,1,\cdots,T_{j}(a)-1$. The last step is by the fact that $\frac{\Delta_{a,j}^{2}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq 1$. For the same reason, we have
$\sqrt{2\pi n\Delta^{2}}\leq$ $\sqrt{\frac{\pi}{50}}$
$\frac{n\Delta^{2}}{2}\leq$ $\frac{1}{200}.$
Not hard to validate that $\sqrt{\frac{1}{50}}+\sqrt{\frac{\pi}{50}}+\frac{1}{200}<0.4$. Applying Lemma 7, we prove the result. ∎
Based on the Lemma 10, we can prove a lower bound connecting $\mathbb{E}N_{A}(\tau)$ and $\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)$, which is the following Lemma.
Lemma 11.
Following the same notation as Lemma 10 and still assume $\delta<\frac{1}{e}$, we have
$\mathbb{E}_{\nu}N_{a}(\tau)\geq\frac{T_{j}(a)}{4}\Pr_{\nu}\left(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right)$
holds for all $j\in[m],a\in[K],a\neq j$.
Proof of Lemma 11.
We firstly prove a claim: Assume $\tau_{n+1}^{j}(a)<\tau$, for all $t$ such that $\tau_{n}+1^{j}(a)<t<\tau_{n+1}^{j}(a)+1$, we have $A_{t}\neq a$. We prove this claim by contradiction. Assume there exists $t$ such that $\tau_{n}^{j}(a)+1<t<\tau_{n+1}^{j}(a)+1$, $A_{t}=a$. Then, from the definition of $\tau_{n}^{j}(a)$, we know $N_{j}(\tau_{n}^{j}(a))+N_{a}(\tau_{n}^{j}(a))=n,N_{j}(\tau_{n}^{j}(a)+1)+N_{a}(\tau_{n}^{j}(a)+1)=n+1$. Since $t>\tau_{n}^{j}(a)+1$, which means $t-1\geq\tau_{n}^{j}(a)+1$, we can conclude
$N_{j}(t)+N_{a}(t)$
$=$ $\mathds{1}(A_{t}=j)+\mathds{1}(A_{t}=a)+N_{j}(t-1)+N_{a}(t-1)$
$=$ $1+N_{j}(t-1)+N_{a}(t-1)$
$\geq$ $1+N_{j}(\tau_{n}^{j}(a)+1)+N_{a}(\tau_{n}^{j}(a)+1)$
$=$ $n+2.$
However, $t\leq\tau_{n+1}^{j}(a)$ suggests that $N_{j}(t)+N_{a}(t)\leq N_{1}(\tau_{n+1}^{j}(a))+N_{a}(\tau_{n+1}^{j}(a))=n+1$, which leads to a contradiction. Then we can rewrite $\mathbb{E}_{\nu}N_{a}(\tau)$ as the following.
$\mathbb{E}_{\nu}N_{a}(\tau)$
$=$ $\mathbb{E}\sum_{t=1}^{+\infty}\mathds{1}(A_{t}=a)\mathds{1}(\tau\geq t)$
$=$ $\mathbb{E}\sum_{n=0}^{+\infty}\mathds{1}(A_{\tau_{n}^{j}(a)+1}=a)\mathds{1}(\tau>\tau_{n}^{j}(a))$
$\geq$ $\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big).$
From Lemma 10, we know
$kl\Bigg(\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big),1-\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)<0.4$
where $kl(x,y)=x\log\frac{x}{y}+(1-y)\log\frac{1-y}{1-x}$. From Pinsker’s Inequality, we can further conclude
$2\Bigg(1-2\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)^{2}<0.4$
$\Rightarrow$ $\frac{1-\sqrt{0.2}}{2}<\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)$
$\Rightarrow$ $\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big)>\frac{1}{4}\Pr_{\nu}\Big(\tau>\tau_{n}^{j}(a)\Big).$
The last line suggests that
$\mathbb{E}_{\nu}N_{a}(\tau)$
$\geq$ $\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big)$
$\geq$ $\frac{1}{4}\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(\tau>\tau_{n}^{j}(a)\Big)$
$\geq$ $\frac{1}{4}\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)$
$=$ $\frac{T_{j}(a)}{4}\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big),$
which is the conclusion we wish to prove. ∎
Given Lemma 11, we are able to take the optimal value of an optimization problem as the lower bound of $\mathbb{E}_{\nu}\tau$, which is Lemma 6.
Proof of Lemma 6.
For any $j\in[m]$, by Lemma 11, we have
$\mathbb{E}\tau$
$=$ $\mathbb{E}N_{j}(\tau)+\sum_{a\neq j}\mathbb{E}\sum_{n=0}^{+\infty}\mathds{1}(A_{\tau_{n}^{j}(a)+1}=a)\mathds{1}(\tau>\tau_{n}^{j}(a))$
$\geq$ $\mathbb{E}N_{j}(\tau)\mathds{1}(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})+\sum_{a\neq j}\mathbb{E}\sum_{n=0}^{T_{j}(a)-1}\mathds{1}(A_{\tau_{n}^{j}(a)+1}=a)\mathds{1}(\tau>\tau_{n}^{j}(a))$
$\geq$ $\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})+\sum_{a\neq j}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)$
$\geq$ $\frac{1}{100\Delta_{j,0}^{2}}\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)+\sum_{a\neq j}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)$
$\geq$ $\sum_{a=1}^{K}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big).$
Thus, we can conclude
$\mathbb{E}\tau$
$\geq$ $\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)$
$\geq$ $\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)}{800\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}\right)}$ (18)
Meanwhile, we can also derive
$\mathbb{E}\tau\geq$ $\sum_{j=1}^{m}\mathbb{E}\tau\mathds{1}(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})$
$\geq$ $\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})$ (19)
Merge (18) and (19), notice that Lemma 9 asserts
$\sum_{j=1}^{m}\Pr_{\nu}\Big(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)\geq\frac{1}{2},$
we know $\mathbb{E}\tau$ must be lower bounded by the optimal value of the following optimization problem.
$\begin{array}[]{cl}\min&\frac{1}{800}\max\left\{\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j},\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\right\}\\ s.t.&\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\end{array}$
Then, we turn to show $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$ holds for $j\in[m],a\in[K]$. If $\mu_{a}\geq\mu_{j}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. If $\mu_{j}>\mu_{a}>\mu_{0}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}=\frac{1}{\Delta_{j,0}^{2}}>\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. If $\mu_{0}\geq\mu_{a}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}=\frac{1}{\Delta_{j,a}^{2}}>\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. Thus, we know $\mathbb{E}\tau$ must be lower bounded by the optimal value of the following optimization problem.
$\begin{array}[]{cl}\min&\frac{1}{3200}\max\left\{\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j},\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\right\}\\ s.t.&\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\end{array}$
Reformulate the above optimization problem, we can turn to solve
$\begin{array}[]{cl}\min&v\\ s.t.&v\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}\\ &v\geq\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)},\forall j\in[m]\\ &\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\\ &v\geq 0.\end{array}$
Given Lemma 6, the last step is to prove a lower bound of the optimal value of (2).
Lemma 12.
Denote opt as the optimal value of problem (2). We can prove
$\text{opt}\geq\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right)$
Proof.
Proof of Lemma 12
We first show function $f(p)=\frac{p}{1+\log\frac{1}{p}}$ is convex for $p\in(0,1]$. Easy to calculate
$\frac{\partial\frac{p}{1+\log\frac{2}{p}}}{\partial p}=$ $\frac{1+\log\frac{2}{p}-p(-\frac{1}{p})}{(1+\log\frac{2}{p})^{2}}$
$=$ $\frac{1}{1+\log\frac{2}{p}}+\frac{1}{(1+\log\frac{2}{p})^{2}}$
$\frac{\partial^{2}\frac{p}{1+\log\frac{2}{p}}}{\partial p^{2}}=$ $\left(-\frac{1}{(1+\log\frac{2}{p})^{2}}-\frac{2}{(1+\log\frac{2}{p})^{3}}\right)\left(-\frac{1}{p}\right)$
$=$ $\left(\frac{1}{(1+\log\frac{2}{p})^{2}}+\frac{2}{(1+\log\frac{2}{p})^{3}}\right)\frac{1}{p}.$
It is straight forward to see $\frac{\partial^{2}\frac{p}{1+\log\frac{2}{p}}}{\partial p^{2}}>0$ holds for all $p\in(0,1]$, suggesting that function $\frac{p}{1+\log\frac{1}{p}}$ is convex, further indicating (2) is a convex optimization problem.
Introduce dual variables $\alpha\geq 0,\gamma\geq 0,\{\beta_{j}\}_{j=1}^{m}\in\mathbb{R}_{+}^{m}$ and Lagrangian function
$L(\alpha,\gamma,\{\beta_{j}\}_{j=1}^{m};\{p_{j}\}_{j=1}^{m},v)$
$=$ $v+\left(\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}-v\right)\alpha+$
$\sum_{j=1}^{m}\left(\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}-v\right)\beta_{j}+(\frac{1}{2}-\sum_{j=1}^{m}p_{j})\gamma$
$=$ $v(1-\alpha-\sum_{j=1}^{m}\beta_{j})+\sum_{j=1}^{m}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}+\beta_{j}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}-\gamma p_{j}\right)+\frac{\gamma}{2}.$
Take $\alpha^{0}=\frac{1}{2}$, $\beta_{j}^{0}=\frac{\frac{1}{2j}}{2\sum_{s=1}^{m}\frac{1}{2s}}$, $\gamma^{0}=\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\frac{1}{2}\cdot+\frac{\frac{1}{2j}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}}{(1+\log(m+1))(1+\log 200+\log m)}$. Since (2) is a convex optimization problem, by the strong duality, we can assert
$\text{opt}=$ $\max_{\alpha\geq 0,\gamma\geq 0,\{\beta_{j}\}_{j=1}^{m}\in\mathbb{R}_{+}^{m}}\min_{\{p_{j}\}_{j=1}^{m}\in[0,1]^{m},v\geq 0}L(\alpha,\gamma,\{\beta_{j}\};\{p_{j}\},v)$
$\geq$ $\min_{\{p_{j}\}_{j=1}^{m}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\};\{p_{j}\},v).$
In the following, we suffice to prove
$\min_{\{p_{j}\}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\}_{j=1}^{m};\{p_{j}\}_{j=1}^{m},v)$
$\geq$ $\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$
Before that, we prove a claim
$\min_{1\geq p_{j}\geq 0}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p_{j}+\frac{p_{j}}{1+\log\frac{2}{p_{j}}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p_{j}\right)\geq-\frac{\gamma^{0}}{100m}$ (20)
Define $\bar{L}_{j}(p)=\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p+\frac{p}{1+\log\frac{2}{p}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p$ and $p^{*}_{j}=\arg\min\limits_{1\geq p\geq 0}\bar{L}_{j}(p)$. Not hard to validate $\bar{L}_{j}(p)$ is a convex function. We prove the claim by discussing two cases.
If $\frac{\partial\bar{L}_{j}(p)}{\partial p}\geq 0$ holds for all $0\leq p\leq 1$, we have
$\min_{1\geq p\geq 0}\bar{L}_{j}(p)=\lim_{p\rightarrow 0}\bar{L}_{j}(p)=0>-\frac{\gamma^{0}}{100m},$
suggesting that the claim (20) is true.
If $\frac{\partial\bar{L}_{j}(p)}{\partial p}<0$ for some $p$, by the fact that $\bar{L}_{j}(p)$ is convex, we can assert there exists $p_{0}\in(0,1)$ such that $\bar{L}_{j}(p)$ is decreasing in the interval $(0,p_{0})$, and $\bar{L}_{j}(p^{*}_{j})<0$. Then, we can prove $p^{*}_{j}<\frac{1}{100m}$ by contradiction. If $p^{*}_{j}\geq\frac{1}{100m}$, we have
$\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p^{*}_{j}+\frac{p^{*}_{j}}{1+\log\frac{2}{p^{*}_{j}}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p^{*}_{j}$
$\geq$ $\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p^{*}_{j}+\frac{p^{*}_{j}}{1+\log 200+\log m}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p^{*}_{j}$
$=$ $p^{*}_{j}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}+\frac{\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}}{1+\log 200+\log m}-\gamma^{0}\right)$
$\geq$ $p^{*}_{j}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\frac{1}{2}+\frac{\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\frac{\frac{1}{2j}}{2\sum_{s=1}^{m}\frac{1}{2s}}}{1+\log 200+\log m}-\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\frac{1}{2}-\frac{\sum_{a=1}^{K}\frac{1}{2j\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}}{(1+\log 200+\log m)(1+\log(m+1))}\right)$
$>$ $0.$
The second last line is by the definition of $\gamma^{0}$. The last line is by the fact that $2\sum_{s=1}^{m}\frac{1}{2s}<1+\log(m+1)$ for $m\geq 1$. This inequality contradicts with the fact that $\bar{L}_{j}(p^{*}_{j})<0$. Thus, we can conclude
$\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p^{*}_{j}+\frac{p^{*}_{j}}{1+\log\frac{2}{p^{*}_{j}}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p^{*}_{j}$
$\geq$ $-\gamma^{0}p^{*}_{j}$
$\geq$ $-\frac{\gamma^{0}}{100m}.$
That means in the case $\frac{\partial\bar{L}_{j}(p)}{\partial p}<0$, we can still conclude the claim (20) is true.
Based on the (20), we have
$\min_{\{p_{j}\}_{j=1}^{m}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\};\{p_{j}\},v)$
$=$ $\min_{v\geq 0}v(1-\alpha^{0}-\sum_{j=1}^{m}\beta_{j}^{0})+$
$\sum_{j=1}^{m}\min_{1\geq p_{j}\geq 0}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}+\beta_{j}^{0}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}-\gamma^{0}p_{j}\right)+\frac{\gamma^{0}}{2}$
$\geq$ $\sum_{j=1}^{m}-\frac{\gamma^{0}}{100m}+\frac{\gamma^{0}}{2}$
$\geq$ $\frac{\gamma^{0}}{4}.$
The second last line is by the fact that $1-\alpha^{0}-\sum_{j=1}^{m}\beta_{j}^{0}=1-\frac{1}{2}-\sum_{j=1}^{m}\frac{\frac{1}{2j}}{2\sum_{s=1}^{m}\frac{1}{2s}}=0$ and the claim (20). From the definition of $\gamma^{0}$, we know
$\min_{\{p_{j}\}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\};\{p_{j}\},v)$
$\geq$ $\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right),$
which also completes the proof. ∎
Appendix C Proof of Upper Bound
C.1 Rigorous Presentation of Algorithm
$U(t,\delta)=\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}.$ (21)
where $\lceil x\rceil:=\max\{x,1\}$
Before presenting the full statement of the algorithm, we first introduce some notations
$N_{a}^{\text{ee}}(\mathcal{H}^{ee})=\sum_{s=1}^{|\mathcal{H}^{ee}|}\mathds{1}(A_{s}(\mathcal{H}^{\text{ee}})=a),N_{a}^{\text{et}}(\mathcal{H}^{et})=\sum_{s=1}^{|\mathcal{H}^{et}|}\mathds{1}(A_{s}(\mathcal{H}^{\text{et}})=a)$ (22)
$\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee})=\frac{\sum_{s=1}^{|\mathcal{H}^{ee}|}\mathds{1}(A_{s}(\mathcal{H}^{\text{ee}})=a)X_{s}(\mathcal{H}^{\text{ee}})}{N_{a}^{\text{ee}}(\mathcal{H}^{ee})},$
$\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})=\frac{\sum_{s=1}^{|\mathcal{H}^{et}|}\mathds{1}(A_{s}(\mathcal{H}^{\text{et}})=a)X_{s}(\mathcal{H}^{\text{et}})}{N_{a}^{\text{et}}(\mathcal{H}^{et})}$
$\text{UCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta)=\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee})+U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
$\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta)=\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee})-C\cdot U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
$\text{UCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)=\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})+U(N_{a}^{\text{et}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
$\text{LCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)=\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})-U(N_{a}^{\text{et}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
where we consider $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$ as stacks that can add index to the stored item, and denote $A_{s}(\mathcal{H}^{\text{ee}}),X_{s}(\mathcal{H}^{\text{ee}})$ as the action and reward stored with index $s$ in $\mathcal{H}^{\text{ee}}$. Same idea applies to $A_{s}(\mathcal{H}^{\text{et}}),X_{s}(\mathcal{H}^{\text{et}})$. If $N_{a}^{\text{ee}}(\mathcal{H}^{ee})=0$ holds for some $a\in[K]$, we take $\text{UCB}^{\text{ee}}_{a}=+\infty,\text{LCB}^{\text{ee}}_{a}=$-$\infty$. Same idea applies to $N_{a}^{\text{et}}(\mathcal{H}^{et}),\text{UCB}^{\text{et}}_{a},\text{LCB}^{\text{et}}_{a}$. Given the above notation, we introduce algorithm 2, which is adapted from Li and Cheung (2025). The following analysis is very similar to the Section 4 and Appendix B in Li and Cheung (2025), but we still remain the full description and proof for completeness. In the following subsection, we discuss each part of the algorithm line by line and highlight our adaptation of history literature in Appendix C.1.5.
1Input: Bracket $B$, Threshold $\mu_{0}$, Tolerance level $\delta$, Tunable parameter $C>1$, Arm number $K$;
2
3Initialize: $\delta_{k}=\frac{1}{3^{k}},\beta_{k}=2^{k},\alpha_{k}=5^{k},\forall k\in\mathbb{N}$, History $\mathcal{H}^{\text{ee}}=\mathcal{H}^{\text{et}}=\emptyset$, Temporary Container $Q=\emptyset$. Introduce notations in (22), $t^{\text{ee}}=t^{\text{et}}=0$;
4 for phase index $k=1,2,\cdots$ do
 /* Phase $k$ starts */
5 while True do
6    if $t^{\text{ee}}\geq 1$ and $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})\geq\mu_{0}$ then
7         Take $\hat{a}_{k}\leftarrow A_{t}^{\text{ee}}$;
8       $\mathcal{H}^{\text{ee}}\leftarrow\mathcal{H}^{\text{ee}}\setminus\{(A^{\text{ee}}_{t},X)\}$, $N_{A^{\text{ee}}_{t}}^{\text{ee}}\leftarrow N_{A^{\text{ee}}_{t}}^{\text{ee}}-1$, $Q\leftarrow Q\cup\{(A^{\text{ee}}_{t},X)\}$, $\hat{a}\leftarrow A_{t}^{\text{ee}}$;
9       while $N_{\hat{a}}^{\text{et}}\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}-1$ do
10            Sample $X\sim\rho_{\hat{a}}$, $\mathcal{H}^{\text{et}}\leftarrow\mathcal{H}^{\text{et}}\cup\{(\hat{a},X)\}$, $N_{\hat{a}}^{\text{et}}\leftarrow N_{\hat{a}}^{\text{et}}+1$, $t^{\text{et}}\leftarrow t^{\text{et}}+1$;
11          if $\text{LCB}^{\text{et}}_{\hat{a}}(\mathcal{H}^{\text{et}},\frac{\delta}{\alpha_{k}})>\mu_{0}$ then
12               Return $\hat{a}$ as a qualified arm;
13
14            end if
15
16         end while
        Break;
       // Enter phase $k+1$
17
18      end if
19    if $\forall a\in B$, $\text{UCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})\leq\mu_{0}$ then
20       if $\delta_{k}\leq\frac{\delta}{3}$ and $|B|=K$ then
21            Take $\hat{a}_{k}=\textsf{None}$ and return None;
          /* Only the bracket with size $K$ is allowed to output None */
22
23       else
            Take $\hat{a}_{k}=\textsf{Not Completed}$ and break;
          // Enter phase $k+1$
24
25         end if
26
27      end if
28    if $N_{a}^{\text{ee}}>(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1,\forall a\in B$ then
         Take $\hat{a}_{k}=\textsf{Not Completed}$ and break;
       // Enter phase $k+1$
29
30      end if
31      Determine $A^{\text{ee}}_{t}=\arg\max\limits_{a\in B,N_{a}^{\text{ee}}\leq(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1}\text{UCB}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})$;
32    if $\exists(a,X)\in Q$ such that $a=A^{\text{ee}}_{t}$ then
33       $\mathcal{H}^{\text{ee}}\leftarrow\mathcal{H}^{\text{ee}}\cup\{(a,X)\}$, $N_{A^{\text{ee}}_{t}}^{\text{ee}}\leftarrow N_{A^{\text{ee}}_{t}}^{\text{ee}}+1$, $Q\leftarrow Q\setminus\{(a,X)\}$;
34
35    else
36         Sample $X\sim\rho_{A_{t}^{\text{ee}}}$, $\mathcal{H}^{\text{ee}}\leftarrow\mathcal{H}^{\text{ee}}\cup\{(A^{\text{ee}}_{t},X)\}$, $N_{A^{\text{ee}}_{t}}^{\text{ee}}\leftarrow N_{A^{\text{ee}}_{t}}^{\text{ee}}+1$, $t^{\text{ee}}\leftarrow t^{\text{ee}}+1$;
37
38      end if
39
40   end while
  Take $\tau^{\text{ee}}_{k}(B)\leftarrow t^{\text{ee}},\tau^{\text{et}}_{k}(B)\leftarrow t^{\text{et}}$;
 // Independent of algorithm design, only for analysis
 /* Phase $k$ ends */
41
42 end for
Algorithm 2 Sequential Exploration Exploitaion (SEE)
C.1.1 From Line 2 to 2: Input and Initialized Parameter
Algorithm 2 takes bracket $B$, threshold $\mu_{0}$, tolerance level $\delta$, arm number $K$, tunable parameter $C$ as input, and only pull arms in $B$. $\mu_{0},K$ are the same as what Algorithm 1 receives, while Algorithm 1 sets $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$ when calling Algorithm 2. We take $C=1.01$ by default, but we can use any value greater than $1$.
Besides input parameter specified by the problem nature and Algorithm 1, 2 also initialize some parameters to manage the pulling process. Parameter $\{\delta_{k}=\frac{1}{3^{k}},\beta_{k}=2^{k},\alpha_{k}=5^{k}\}_{k=1}^{+\infty}$ are used to set up the tolerance level and budget during the pulling process. It is possible to specify other values for $\delta_{k},\beta_{k},\alpha_{k}$, see Appendix B.1 in Li and Cheung (2025). $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}},Q$ are used to store reward collected from pulling arms. As we have discussed above, $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$ are two stacks that can maintain the arrival order of reward of each arm $a\in B$. This requirement is to guarantee we are able to transfer the latest collected reward out of $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$. $Q$ is used to guarantee $\text{LCB}_{a}^{\text{ee}}$ cannot be greater than $\mu_{0}$ at the start of each phase, which will be discussed in section C.1.4.
C.1.2 Line 2 to Line 2: Overall Idea of Algorithm 2
At the start of Algorithm 2, it splits the arm distributions into two parallel and independent copies, one for exploration period, and another for exploitation period. Correspondingly, reward collected in the exploration periods are stored in $\mathcal{H}^{\text{ee}}$. Reward collected in the exploration periods, which is from Line 2 to Line 2, are stored in $\mathcal{H}^{\text{et}}$.
After specifying the parameters through Line 2 and 2, Algorithm 2 splits the pulling process into multiple phases, noted by the For loop in Line 2. In each phase $k$, Algorithm 2 consists of two parts. The first part is the conditions from Line 2 to Line 2. The second part, which is called Exploration Period, ranges from Line 2 to 2. Exploration Period adopts UCB rule in Jamieson and Nowak (2014) to pull arms. These two parts are sequentially executed in the While Loop at Line 2. Though the condition of the While Loop is always True, but the conditions in the first part is able to guarantee that the Algorithm 2 never gets stuck in a infinity loop. See the discussion in section C.1.3.
At a phase $k$, the exploration period utilizes $\delta_{k}$ as the tolerance level to apply UCB rule of pulling arms, while the exploitation period utilizes $\frac{\delta}{\alpha_{k}}$ as the tolerance level to calculate Upper and Lower Confidence Bounds and only pull a single arm. Once the tolerance level of both periods are below $\frac{\delta}{3}$ and $\hat{a}$ takes value in $K]\cup\{\textsf{None}\}$, Algorithm 2 will terminate. What’s more, Algorithm 2 also sets up a maximum pulling times for both periods. For the exploration period at phase $k$, the maximum pulling times for each arm $a\in B$ is $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$. For the exploitation period at phase $k$, the maximum pulling times of the single arm is $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$. Section C.1.3 and C.1.4 provide more details.
C.1.3 Line 2 to 2: Conditions on Phase Transition
There are three possible conditions to end the exploration period in a phase $k$. The first one is Line 2, judge whether the algorithm should enter the exploitation period. Intuitively speaking, the inequality $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})\geq\mu_{0}$ provides a recommended qualified arm with confidence $1-\Theta(\delta_{k})$. Since $A_{t}^{\text{ee}}$ is the action we take in the last round, this condition also guarantees $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ holds before the collecting the latest sample of $A_{t}^{\text{ee}}$. To further confirm whether arm $A_{t}^{\text{ee}}$ is qualified, we need to keep pulling arm $A_{t}^{\text{ee}}$ by collecting samples independent of the exploration periods. Line 2 suggests that the algorithm will terminate and output $A_{t}^{\text{ee}}$, if the LCB calculated by the tolerance level $\delta/\alpha_{k}$ is above $\mu_{0}$ during the exploitation process. Before entering the while loop (Line 2) in the exploitation period, Line 2 transfer the latest collected sample of $A_{t}^{\text{ee}}$ in the exploration period into $Q$. Usage of $Q$ is to guarantee Lemma 22, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. Similar to the reason described in the section 4.1 of Li and Cheung (2025), Lemma 22 is required to prove Lemma 23, further Theorem 3. We put more discussion of $Q$ to section C.1.4.
The second condition is Line 2, aiming to output None for negative case. To be specific, if the Upper Confidence Bound(UCB) in the exploration periods are all below $\mu_{0}$, Algorithm 2 will leave the exploration periods in phase $k$. Then, the condition in Line 2 determines the next period. If the bracket $B$ is indeed $[K]$ and the tolerance level $\delta_{k}<\frac{\delta}{3}$, Algorithm 2 will terminate the overall pulling process and output None. In other cases, Algorithm 2 will enter the exploration period in the next phase $k+1$. In the case $\max_{a\in B}\mu_{a}<\mu_{0}$ and $|B|<K$, it is possible that Algorithm 2 never break the for loop in Line 2.
The third condition is Line 2, which is correlated to the budget of the exploration period. For a phase $k$, Algorithm 2 sets up a maximum pulling times $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ for each arm $a\in B$. Recall we only consider picking $A^{\text{ee}}_{t}$ from those arms whose total pulling times from phase 1 to phase $k$ is below $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ in Line 2. The third condition is Line 2 is to enter the exploration period in phase $k+1$, if pulling times are no less than $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1$ holds for all $a\in B$. Line 2 and 2 also guarantee Algorithm 2 will not get stuck into a infinite loop of “While True” in Line 2.
C.1.4 Line 2 to 2: Exploration
In phase $k$, Algorithm 2 use tolerance level $\delta_{k}$ to calculate Upper Confidence Bound(UCB) according to (22) and determine the arm $A_{t}^{\text{ee}}$ by Line 2. Line 2 to 2 determine where to collect sample of arm $A_{t}^{\text{ee}}$. If $Q$ contains one sample of arm $A_{t}^{\text{ee}}$, Algorithm 2 will transfer this sample to $\mathcal{H}^{\text{ee}}$ without collecting a new sample from distribution $\nu_{A_{t}^{\text{ee}}}$. If not, Algorithm 2 samples a new random variable from $\nu_{A_{t}^{\text{ee}}}$.
As we mentioned in section C.1.3, usage of $Q$ is to guarantee Lemma 22, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. Someone may question whether we can directly abandon the latest collected sample of $A_{t}^{\text{ee}}$ in $\mathcal{H}$, which seems also able to guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. We cannot to do so. By transferring samples into or out $Q$, we are able to derive
Lemma 13.
Given an algorithm copy alg taking bracket $B$ as input. Condition on reward sequence in the exploration period $\{X_{a,s}^{\text{ee}}\}_{s=1}^{+\infty}$, for each $a\in B$, we can assert $\hat{\mu}_{a}(\mathcal{H}^{\text{ee}})\in\{\frac{\sum_{s=1}^{t}X_{a,s}^{\text{ee}}}{t}\}_{t=1}^{+\infty}$ throughout the pulling process.
The reason is we do not change the arrival order of $\{X_{a,s}^{\text{ee}}\}_{s=1}^{+\infty}$ for each $a\in B$. To analyze Algorithm 2, we need to introduce
$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\right\}$
$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\right\}.$
where $\lceil x\rceil^{+}=\max\{x,1\}$. $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately. More details are in section C.2. Lemma 13 suggests that for phase $k\geq\kappa^{\text{ee}}_{b}$, we can guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{a}<\text{UCB}_{a}^{\text{ee}}$ holds for all $a\in B_{b}$, which further helps to prove upper bounds of the pulling times for all $a\in B_{b}$. Simply dropping the latest collected samples are unable to guarantee 13, making it much harder to conduct the analysis.
Besides, we also care about the size of $Q$ during the pulling process, as we can observe the following Lemmas.
Lemma 14.
Given an algorithm copy alg, throughout the pulling process, we have
$t^{\text{ee}}=|\mathcal{H}^{\text{ee}}|+|Q|$
always hold.
The reason is all the collected samples in the exploration periods are either in $\mathcal{H}^{\text{ee}}$ or $Q$. Notice that we only transfer sample from $\mathcal{H}^{\text{ee}}$ to $Q$ in Line 2. In this case, $A_{t}^{\text{ee}}$ just get pulled in the last round of exploration. Meanwhile, once $Q$ contains a sample for an arm $a\in B$, condition in 2 guarantees Algorithm 2 will never collect new samples of arm $a$ in the exploration period. This property implies following.
Lemma 15.
Consider an algorithm copy alg taking bracket $B$ as input, we have $|Q|\leq|B|$ throughout the pulling process
At the end of this subsection, we present a Lemma similar to 13, but for the exploitation period.
Lemma 16.
Given an algorithm copy alg taking bracket $B$ as input. Condition on $\{X_{a,s}^{\text{et}}\}_{s=1}^{+\infty}$, for each $a\in B$, we can assert $\hat{\mu}_{a}(\mathcal{H}^{\text{et}})\in\{\frac{\sum_{s=1}^{t}X_{a,s}^{\text{et}}}{t}\}_{t=1}^{+\infty}$ throughout the pulling process.
This is correct, as we collect the samples in the exploitation period sequentially in $\mathcal{H}^{\text{et}}$.
C.1.5 Adaptation of History Literature
Compared to the Algorithm SEE in Li and Cheung (2025), our main adaptation consists of two points. The first adaptation is to put the stopping conditions (Line 2 to 2) before the exploration periods (Line 2 to 2). This design is to guarantee Lemma 29, guaranteeing Algorithm 2 will not pull arm $a$ whose $\text{UCB}_{a}^{\text{ee}}$ is below $\mu_{0}$.
The second point is for each phase $k$, we set up a maximum pulling times of each arm in the exploration period, instead of maximum pulling times of the total pulling times (reference line index is to be added). The reason is the arm in a bracket $B$ is randomized, it is hard to analyze the required pulling times that a bracket need to output a $\hat{a}_{k}$.
C.2 Properties of Algorithm 2
This subsections follows the Appendix B in Li and Cheung (2025), which are nearly the same. Here we remain the full proof for completeness, as we have modified some implementation details and proof.
In Algorithm 1, we create $\lceil\log_{2}K\rceil+1$ independent copies of Algorithm 2. For each copy index $b\in[\lceil\log_{2}K\rceil+1]$, we further create two independent sampling copies for exploration and exploitation periods separately. We define
$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\right\}$
$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\right\}.$
where $\lceil x\rceil^{+}=\max\{x,1\}$. $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately. The above definition of $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ are conditioned on the realized reward of $\{X_{a,s}^{\text{ee}}\}_{a=1,s=1}^{K,+\infty}$, $\{X_{a,s}^{\text{et}}\}_{a=1,s=1}^{K,+\infty}$ collected by the algorithm copy $b$. $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ are the minimum phase index such that concentration event holds. Since $B_{b}$ is determined by an exogenous random permutation, we know $B_{b}$ is independent with $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$.
In the following, we will analyze the properties of Algorithm 2, a given bracket $B$. For convenience, given a bracket $B$ with index $b$, we denote $\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B}$ as corresponding $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$, without emphasizing the index $b$. Further, we denote $\tau_{k}^{\text{ee}}(B),\tau_{k}^{\text{et}}(B)$ as the value of $t^{\text{ee}},t^{\text{et}}$ at the end of phase $k$, i.e. each time we enter Line 2 in Algorithm 2.
Lemma 17.
For any $B$, we have $\Pr(\kappa^{\text{ee}}_{B}\geq k)\leq\frac{\pi^{2}}{6}\delta_{k-1}$, $\Pr(\kappa^{\text{ee}}_{B}\geq k)\leq\frac{\pi^{2}}{6}\frac{\delta}{\alpha_{k-1}}$
Proof of Lemma 17.
By the Lemma 35, we have
$\Pr(\kappa^{\text{ee}}_{B}\geq k)\leq$ $\Pr\left(\exists t,\exists a\in[K],|\hat{\mu}^{ee}_{a}(t)-r^{\nu}_{a}|\geq\sqrt{\frac{4\log\frac{2K}{\delta_{k-1}}+8\log\log(2t)}{t}}\right)$
$\leq$ $\sum_{a=1}^{K}\Pr\left(\exists t,|\hat{\mu}^{ee}_{a}(t)-r^{\nu}_{a}|\geq\sqrt{\frac{4\log\frac{2K}{\delta_{k-1}}+8\log\log(2t)}{t}}\right)$
$\leq$ $\sum_{a=1}^{K}\frac{\pi^{2}\delta_{k-1}}{6K}$
$=$ $\frac{\pi^{2}\delta_{k-1}}{6}$
and
$\Pr(\kappa^{\text{et}}_{B}\geq k)\leq$ $\Pr\left(\exists t,\exists a\in[K],|\hat{\mu}^{et}_{a}(t)-r^{\nu}_{a}|\geq\sqrt{\frac{4\log\frac{2K\alpha_{k-1}}{\delta}+8\log\log(2t)}{t}}\right)$
$\leq$ $\sum_{a=1}^{K}\Pr\left(\exists t,|\hat{\mu}^{ee}_{a}(t)-r^{\nu}_{a}|\geq\sqrt{\frac{4\log\frac{2K\alpha_{k-1}}{\delta}+8\log\log(2t)}{t}}\right)$
$\leq$ $\sum_{a=1}^{K}\frac{\pi^{2}\delta}{6K}$
$=$ $\frac{\pi^{2}\delta}{6\alpha_{k-1}}.$
Lemma 18.
Consider an Algorithm copy $b$ with bracket $B$, we have $\Pr(\text{alg copy }b\text{ outputs }\hat{a}\text{ such that }\mu_{\hat{a}}\leq\mu_{0})\leq\frac{\pi^{2}}{6}\frac{\delta}{5}$.
Proof of Lemma 18.
By the Line 2 to 2, we know
$\Pr(\text{alg copy }b\text{ outputs }\hat{a}\text{ such that }\mu_{\hat{a}}<\mu_{0})$
$\leq$ $\Pr\left(\exists a,s.t.\mu_{a}<\mu_{0},t\in\mathbb{N},\hat{\mu}_{a}^{\text{et}}(t)-\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{1}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\geq\mu_{0}\right)$
$\leq$ $\sum_{a:\mu_{a}<\mu_{0}}\Pr\left(t\in\mathbb{N},\hat{\mu}_{a}^{\text{et}}(t)-\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{1}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\geq\mu_{a}\right)$
$\leq$ $\sum_{a:\mu_{a}\leq\mu_{0}}\frac{\pi^{2}}{6}\frac{\delta}{\alpha_{1}K}$
$\leq$ $\frac{\pi^{2}}{6}\frac{\delta}{5}.$
The second last line is by the Lemma 35. ∎
Lemma 19.
Assume $\mu_{1}>\mu_{0}$. Consider an Algorithm copy $\lceil\log_{2}K\rceil+1$, whose allocated bracket is $[K]$. We have $\Pr(\text{alg copy }\lceil\log_{2}K\rceil+1\text{ outputs }\textsf{None})\leq\frac{\pi^{2}}{6}\frac{\delta}{3}$.
Proof of Lemma 19.
By the condition in Line 2, we know
$\Pr(\text{alg copy }\lceil\log_{2}K\rceil+1\text{ outputs }\textsf{None})$
$\leq$ $\Pr(\exists t\in\mathbb{N},\delta_{k}\leq\frac{\delta}{3},\hat{\mu}_{1}^{\text{ee}}(t)+\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\leq\mu_{0})$
$\leq$ $\Pr(\exists t\in\mathbb{N},\hat{\mu}_{1}^{\text{ee}}(t)+\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta/3}}}{t}\leq\mu_{0})$
$\leq$ $\frac{\pi^{2}}{6}\frac{\delta}{3}.$
The last line is by the Lemma 35. ∎
Lemma 20.
For any $B$, and any $a\in B$, we have $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\leq(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$.
Proof of Lemma 20.
We use induction on $k\in\mathbb{N}$. For $k=1$, by the Line 2 in Algorithm 2, we konw
$N_{a}^{\text{ee},B}\Big(\tau_{1}^{\text{ee}}(B)\Big)\leq(C+1)^{2}\beta_{1}\log\frac{4K}{\delta_{1}}.$
If the conclusion holds for $k$, then for $k+1$, we can still conclude
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}\right\}$
$\leq$ $\max\left\{(C+1)^{2}\beta_{k-}\log\frac{4K}{\delta_{k-1}},(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}\right\}$
$=$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}.$
The first step is still caused by the Line 2 in Algorithm 2. The second step is from the foundation. The last line shows that the induction holds. ∎
Lemma 21.
For any $B$, $\tau_{k}^{\text{ee}}(B)\leq|B|(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$, $\tau_{k}^{\text{et}}(B)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$
Proof of Lemma 21.
Regarding $\tau_{k}^{\text{ee}}(B)\leq|B|(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$, we can directly conclude this from Lemma 20.
Regarding $\tau_{k}^{\text{et}}(B)\leq\sum_{k^{\prime}=1}^{k}\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k^{\prime}}\log\frac{4K\alpha_{k^{\prime}}}{\delta}$, we suffice to show $\tau_{k}^{\text{et}}(B)-\tau_{k-1}^{\text{et}}(B)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$ holds for any phase index $k$. By direct calculation, we have
$\tau_{k}^{\text{et}}(B)-\tau_{k-1}^{\text{et}}(B)$
$=$ $\sum_{a=1}^{K}N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)-N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k-1}^{\text{et}}(B)\Big)\mathds{1}(\hat{a}_{k}=a)$
$\leq$ $\sum_{a=1}^{K}N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)\mathds{1}(\hat{a}_{k}=a)$
$\leq$ $\sum_{a=1}^{K}\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}\mathds{1}(\hat{a}_{k}=a)$
$\leq$ $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}.$
The second last line is by the loop condition in Line 2 of Algorithm 2. ∎
Lemma 22.
For any $B$, at the start of each phase $k$, i.e. each time we enter Line 2 in Algorithm 2, we have $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ holds for all $a\in B$.
Proof of Lemma 22.
We apply induction on $k$. For $k=1$, this is obviously true, as we denote the $\text{LCB}_{a}^{\text{ee}}$ as -$\infty$ if the pulling times of arm $a$ is 0.
If the conclusion holds for the phase k1k-1, we prove the case of $k$ by discussing all the possible values of $\hat{a}_{k-1}$. Before the start of phase $k$, there are three possible scenarios for the value of $\hat{a}_{k-1}$. Phase k1k-1 may end with taking $\hat{a}_{k-1}$ takes value at the Line 2, 2 or 2.
If $\hat{a}_{k-1}$ takes Not Completed at the Line 2, we have we $\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})<\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$, forall $a\in[K]$.
If $\hat{a}_{k-1}$ takes Not Completed at the Line 2, We can also assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$. Prove by contradiction. Assume there exists $a\in[K]$ such that $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})\geq\mu_{0}$ at the end of phase k1k-1. From the induction, since $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$ at the start of phase k1k-1, the algorithm would enter Line 2 and $\hat{a}_{k-1}$ cannot take value Not Completed. Thus, we can assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$ holds at the end of phase k1k-1. Further, at the start of phase $k$, we have $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$.
The remaining work is to prove the conclusion in the case that $\hat{a}_{k-1}\in[K]$ at the Line 2. We can first assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\neq\hat{a}_{k-1}$. The reason is from the induction, $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$ holds for all $a$ at the start of phase k1k-1. Once there exists an arm $a$ whose LCB is above $\mu_{0}$ after an arm pull, the algorithm will enter Line 2, suggesting that there is at most one arm whose LCB is above $\mu_{0}$.
Then, we turn to analyze $\text{LCB}_{\hat{a}_{k-1}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})$ at the start of phase $k$. Before the execution of Line 2, the following inequalities must hold
$\frac{X+\sum_{s=1}^{N_{\hat{a}_{k-1}}^{\text{ee}}-1}X_{\hat{a}_{k-1},s}}{N_{\hat{a}_{k-1}}^{\text{ee}}}-C\cdot U\left(N_{\hat{a}_{k-1}}^{\text{ee}},\frac{\delta_{k-1}}{K}\right)>\mu_{0}$
$\frac{\sum_{s=1}^{N_{\hat{a}_{k-1}}^{\text{ee}}-1}X_{\hat{a}_{k-1},s}}{N_{\hat{a}_{k-1}}^{\text{ee}}-1}-C\cdot U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k-1}}{K}\right)\leq\mu_{0}$
Since we put the last collected sample $X$ into $Q$ in Line 2, we can assert $\hat{\mu}_{\hat{a}_{k-1}}$ before the start of phase $k$ must be $\frac{\sum_{s=1}^{N_{\hat{a}_{k-1}}^{\text{ee}}-1}X_{\hat{a}_{k-1},s}}{N_{\hat{a}_{k-1}}^{\text{ee}}-1}$. Since $U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k}}{K}\right)>U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k-1}}{K}\right)$, we know before the start of phase $k$, $\text{LCB}_{\hat{a}_{k-1}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ must hold.
Following lemma suggests that the exploration period is able to output correct answer, given some concentration event.
Lemma 23.
For $B$ such that $\max_{a\in B}\mu_{a}\geq\mu_{j}$ for some $\mu_{j}>\mu_{0}$, and phase index $k\geq\kappa^{\text{ee}}_{B}$, Algorithm 2 either takes $\hat{a}_{k}\in[K]$ such that $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$ or takes $\hat{a}_{k}=\textsf{Not Completed}$
Proof of Lemma 23.
From the assumption $k\geq\kappa^{\text{ee}}_{B}$, by the Lemma 13, we know
$\hat{\mu}_{a}(\mathcal{H}^{\text{ee}})-U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})\leq\mu_{a}\leq\hat{\mu}_{a}(\mathcal{H}^{\text{ee}})+U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$ (23)
holds for $a\in B$ through out the phase $k$.
We first show that Algorithm 2 never enters Line 2. Denote $a^{*}(B)=\arg\max_{a\in B}\mu_{a}$. From the assumption, we know $\mu_{a^{*}(B)}>\mu_{0}$. Together with the definition $\text{UCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta_{k})=\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee})+U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$, we know $\text{UCB}^{\text{ee}}_{a^{*}(B)}(\mathcal{H}^{ee},\delta_{k})>\mu_{0}$ holds through out the phase $k$, suggesting that the condition in Line 2 never holds. Meanwhile, to quit the While loop in the Line 2, Algorithm 2 must enter one of the Lines in 2, 2, 2. Then, we can conclude $\hat{a}_{k}$ only takes value in $K]\cup\{\textsf{Not Completed}\}$.
The remaining work is to prove if $\hat{a}_{k}\in[K]$, we have $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$. For simplicity, we denote $\tau_{k}^{\text{ee}}$ as $\tau_{k}^{\text{ee}}(B)$, and prove by contradiction. If $\hat{a}\in[K]$ and $\mu_{\hat{a}}<\omega\mu_{j}+(1-\omega)\mu_{0}<\omega\mu_{a^{*}(B)}+(1-\omega)\mu_{0}$, we have
$A_{\tau^{\text{ee}}_{k}}^{\text{ee}}=\hat{a}_{k},\hat{a}_{k}=\arg\max_{i\in B}\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}+U(N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})$
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1,\frac{\delta_{k}}{K})<\mu_{0}.$
where $\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}$ denotes the empirical mean value of arm i of the first $N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)$ samples. The third line is by the Lemma 22, suggesting $\text{LCB}_{\hat{a}_{k}}^{\text{ee}}<\mu_{0}$ must hold before the latest collection of the sample.
As we take $U(t,\frac{\delta}{K})=\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}t\rceil,1\}}\log\frac{2K(\lceil\log_{2}t\rceil)^{2}}{\delta}}}{t}$, we get
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
$\Leftrightarrow$ $\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}>\mu_{0}$
[…]
$\Rightarrow$ $\omega\mu_{a^{*}(B)}+(1-\omega)\mu_{0}-(C-1)\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}>\mu_{0}$
$\Leftrightarrow$ $\omega(\mu_{a^{*}(B)}-\mu_{0})>(C-1)\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}$
$\Leftrightarrow$ $\frac{2\omega(\mu_{a^{*}(B)}-\mu_{0})}{C-1}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$
By $\omega=\frac{C-1}{C+3}$, we have $\frac{(1-\omega)}{2}=\frac{1-\frac{C-1}{C+3}}{2}=\frac{4}{2(C+3)}=\frac{2\omega}{C-1}$, which means
$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$
Notice that
$\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}$
$\leq$ $\frac{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}$
$\leq$ $\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})},$
we can conclude
$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}\geq\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}.$
On the other hands,
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}+U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
[…]
$\Leftarrow$ $\omega\mu_{1}+(1-\omega)\mu_{0}+2U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
$\Leftrightarrow$ $\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})$
$\Leftrightarrow$ $\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}.$
In summary, we can conclude
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau)}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
$\Rightarrow$ $\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}(\tau^{\text{ee}}-1)}+U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
[…]
$\Rightarrow$ $A_{\tau^{\text{ee}}_{k}}^{\text{ee}}\neq\hat{a}_{k}.$
The last line suggests a contradiction. We can conclude $\mu_{\hat{a}}\geq\frac{C-1}{C+3}\mu_{j}+\frac{4}{C+3}\mu_{0}$. ∎
Lemma 24.
For $B$ such that $\max_{a^{\prime}\in B}\mu_{a^{\prime}}\geq\mu_{j}$ for some $\mu_{j}>\mu_{0}$, phase index $k\geq\kappa^{\text{ee}}_{B}$, for each arm $a\in B$, we have
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right\}$
Proof of Lemma 24.
We first prove an intermediate conclusion, which is for each arm $a\in B$, we have
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$ (24)
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right\}$
hols for for any phase index $k\geq\kappa^{\text{ee}}_{B}$. If the above conclusion holds, we can use induction to conclude
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\Bigg\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k-1}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$
$\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\Bigg\}$
$=$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right\},$
which proves the Lemma. Thus, we suffice to prove the intermediate conclusion (24).
From the assumption $k\geq\kappa^{\text{ee}}_{B}$, we know (23) still holds. Fix the arm $a\in B$, we prove the claim by discussing the value of $\mu_{a}$. We consider two cases, $\mu_{a}\geq\frac{\mu_{j}+\mu_{0}}{2}$ or $\mu_{a}<\frac{\mu_{j}+\mu_{0}}{2}$.
For arm $a$ such that $\mu_{a}\geq\frac{\mu_{j}+\mu_{0}}{2}$ holds, we suffice to prove
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{1}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\right\}.$
We can conduct direct calculation, as
$t\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
[…]
$\Leftrightarrow$ $\mu_{a}-(C+1)\sqrt{\frac{4\log\frac{2K(\log_{2}(2t))^{2}}{\delta_{k}}}{t}}>\mu_{0}$
$\Rightarrow$ $\mu_{a}-(C+1)\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}>\mu_{0}$
[…]
By the Lemma 22, we know at the start of phase $k$, $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$, suggesting that $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)\leq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$. Then, we can assert
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)<\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1.$
If not, the algorithm will output $\hat{a}_{k}=a$ before $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$ holds.
For arm $a$ such that $\mu_{a}<\frac{\mu_{j}+\mu_{0}}{2}$ holds, we suffice to prove
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}\right\}.$
We can also conduct direct calculation, with
$t\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}$
[…]
$\stackrel{{\scriptstyle C>1}}{{\Rightarrow}}$ $\mu_{a}+2\sqrt{\frac{4\log\frac{2K(\log_{2}(2t))^{2}}{\delta_{k}}}{t}}\leq\mu_{j}$
$\Rightarrow$ $\mu_{a}+2\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\leq\mu_{j}$
[…]
If $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)\geq\frac{28(C+1)^{2}\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}+1$, we know $\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})<\mu_{j}\leq\mu_{a^{*}(B)}\leq\text{UCB}_{a^{*}(B)}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})$ holds through out the phase $k$, and arm $a$ will never get pulled. We can assert $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)=N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)$.
If $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)<\frac{28(C+1)^{2}\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}+1$, we can assert $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\leq\frac{29(C+1)^{2}\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}$ as the arm $a$ never gets pulled once $\text{UCB}_{a}^{\text{ee}}\leq\mu_{a^{*}(B)}\leq\text{UCB}_{a^{*}(B)}^{\text{ee}}$ holds. We complete the proof. ∎
Lemma 25.
For $B$ such that $\max_{a^{\prime}\in B}\mu_{a^{\prime}}\geq\mu_{j}$, and phase index $k\geq\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{112(C+1)^{2}}{\Delta_{j,0}^{2}}\rceil\}$, Algorithm 2 will take $\hat{a}_{k}\in[K]$ such that $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$.
Proof of Lemma 25.
From the proof of Lemma 23, we know Algorithm 2 will never meet the condition in Line 2.
Meanwhile, from the calculation in the proof of Lemma 24, we know
$t\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a^{*}(B),0}^{2}}\right)}{\Delta_{a^{*}(B),0}^{2}}$
$\Rightarrow$ $\hat{\mu}_{a^{*}(B),t}-C\cdot\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}>\mu_{0}$
and
$k\geq\lceil\log_{2}\frac{112(C+1)^{2}}{\Delta_{j,0}^{2}}\rceil$
$\Rightarrow$ $\beta_{k}\geq\frac{112(C+1)^{2}}{\Delta_{j,0}^{2}},\log\frac{1}{\delta_{k}}\geq\log 3\log_{2}\frac{112(C+1)^{2}}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $\frac{1}{2}\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{56\log\frac{2K}{\delta_{k}}}{\Delta_{j,0}^{2}},\frac{1}{2}\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{56\log\log\frac{24(C+1)^{2}}{\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{56\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{j,0}^{2}}\right)}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{j,0}^{2}}\right)}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a^{*}(B),0}^{2}}\right)}{\Delta_{a^{*}(B),0}^{2}},$
which means the Algorithm will first fulfill the condition in Line 2 before fulfilling the condition in Line 2. This observation concludes $\hat{a}_{k}\in[K]$. By the Lemma 23, we can further conclude $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$. ∎
Lemma 26.
For $B$ such that $\max_{a^{\prime}\in B}\mu_{a^{\prime}}\geq\mu_{j}$, and phase index $k\geq\max\{\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$. If Algorithm 2 enters Line 2 with $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$, Algorithm 2 will terminate and output $\hat{a}_{k}$.
Proof of Lemma 26.
From the condition $k\geq\kappa^{\text{et}}_{B}$, by the Lemma 16, we know
$\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{\text{et}})-U(N_{a}^{\text{et}}(\mathcal{H}^{et}),\frac{\delta}{K\alpha_{k}})\leq\mu_{a}\leq\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{\text{et}})+U(N_{a}^{\text{et}}(\mathcal{H}^{et}),\frac{\delta}{K\alpha_{k}})$
holds for $a\in B$ through out the phase $k$.
By similar calculation in the proof of Lemma 24, we know
$t\geq\frac{112\left(\log\frac{2K\alpha_{k}}{\delta}+\log\log\frac{96}{\omega^{2}\Delta_{j,0}^{2}}\right)}{\omega^{2}\Delta_{j,0}^{2}}$
$\Rightarrow$ $t\geq\frac{112\left(\log\frac{2K\alpha_{k}}{\delta}+\log\log\frac{96}{\Delta_{\hat{a}_{k},0}^{2}}\right)}{\Delta_{\hat{a}_{k},0}^{2}}$
$\Rightarrow$ $\hat{\mu}_{\hat{a}_{k},t}^{\text{et}}-\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}>\mu_{0}.$
Here $\hat{\mu}_{\hat{a}_{k},t}^{\text{et}}$ denotes the empirical mean of first $t$ samples collected in the exploitation period. Notice that
$k\geq\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil$
$\Rightarrow$ $\beta_{k}\geq\frac{448}{\Delta_{j,0}^{2}},\log\alpha_{k}\geq\log\frac{96(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\log 5$
$\Rightarrow$ $\beta_{k}\log\frac{4K\alpha_{k}}{\delta}\geq\frac{448\log\frac{2K\alpha_{k}}{\delta}}{\Delta_{j,0}^{2}},\beta_{k}\log\frac{4K\alpha_{k}}{\delta}\geq\frac{448\log\log\frac{96}{\omega^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $\beta_{k}\log\frac{4K\alpha_{k}}{\delta}\geq\frac{224\left(\log\frac{2K\alpha_{k}}{\delta}+\log\log\frac{96}{\omega^{2}\Delta_{j,0}^{2}}\right)}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}-1\geq\frac{112\left(\log\frac{2K\alpha_{k}}{\delta}+\log\log\frac{96}{\omega^{2}\Delta_{j,0}^{2}}\right)}{\omega^{2}\Delta_{j,0}^{2}}.$
The last line suggests that Algorithm 2 would output $\hat{a}_{k}$ before meeting the stop condition $N_{\hat{a}}^{\text{et}}>\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}-1$ in Line 2. ∎
Lemma 27.
For $B$ such that $\max_{a^{\prime}\in B}\mu_{a^{\prime}}\geq\mu_{j}$, Algorithm 2 must terminate no later than the end of phase $\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$.
Proof of Lemma 27.
This is a direct conclusion from Lemma 25, 26. ∎
Lemma 28.
Consider the largest bracket $B$, i.e. $B=[K]$. If $\max_{a^{\prime}\in B}\mu_{a^{\prime}}=\max_{a^{\prime}\in[K]}\mu_{a^{\prime}}<\mu_{0}$, then for phase index $k\geq\kappa^{\text{ee}}_{B}$, Algorithm 2 either takes $\hat{a}_{k}=\textsf{None}$ or takes $\hat{a}_{k}=\textsf{Not Completed}$
Proof of Lemma 28.
From the assumption $k\geq\kappa^{\text{ee}}_{B}$, by the Lemma 13, we know
$\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}})-U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})\leq\mu_{a}\leq\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}})+U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$ (25)
holds for $a\in B$ through out the phase $k$.
The analysis is similar to the Lemma 23. To quit the While loop in the Line 2, Algorithm 2 must enter one of the Lines in 2, 2, 2. By the assumption $\mu_{a}<\mu_{0}$ holds for all $a\in[K]$ and (28), we know
$\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})$
$=$ $\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}})-C\cdot U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$
$\stackrel{{\scriptstyle C>1}}{{\leq}}$ $\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}})-U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$
$\leq$ $\mu_{a}$
$<$ $\mu_{0},$
holds through out the phase $k$. The last line also suggests that the condition in Line 2 will never be satisfied in phase $k$. Thus, we can conclude $\hat{a}_{k}=\textsf{None}\text{ or }\textsf{Not Completed}$. ∎
Lemma 29.
For $B$ such that $|B|=K$ and $\max_{a^{\prime}\in B}\mu_{a^{\prime}}<\mu_{0}$, phase index $k\geq\kappa^{\text{ee}}_{B}$, for each arm $a\in B$, we have
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\cdot\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\right\}$
Proof of Lemma 29.
Similar to the idea of Lemma 24, we first turn to prove an intermediate conclusion
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$ (26)
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\right\}$
holds for all $k\geq\kappa^{\text{ee}}_{B}$ and $a\in B$. If the above conclusion holds, we can use induction to conclude
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\Bigg\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{113\left(\log\frac{4K}{\delta_{k-1}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}},$
$\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\Bigg\}$
$=$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\right\}.$
which proves the Lemma. Thus, we suffice to prove the intermediate conclusion (26).
We can conduct direct calculation for each arm $a$.
$t\geq\frac{28\cdot 2^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
[…]
$\Leftrightarrow$ $\mu_{a}+2\sqrt{\frac{4\log\frac{2K(\log_{2}(2t))^{2}}{\delta_{k}}}{t}}\leq\mu_{0}$
$\Rightarrow$ $\mu_{a}+2\cdot\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}>\mu_{0}$
[…]
The last line is by the assumption $k\geq\kappa^{\text{ee}}_{B}$, which means (23) still holds.
If $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)\geq\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$, we know $\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})<\mu_{0}$ holds at the start of phase $k$. If $a=\arg\max_{a^{\prime}\in B}\text{UCB}_{a^{\prime}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})$ holds during the phase $k$, Algorithm 2 will turn to phase $k+1$ as the condition in the Line 2 holds. Thus, we can conclude Algorithm 2 will never pull arm $a$ during the phase $k$. That means if $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)\geq\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$, $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)=N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)$.
If $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)<\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$. We can also assert $\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})\leq\mu_{0}$ will hold if $N_{a}^{\text{ee},B}\geq\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$ holds in the futuer rounds. Similar to the above argument, if $\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})\leq\mu_{0}$, $a=\arg\max_{a^{\prime}}\text{UCB}_{a^{\prime}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})$ both holds, the algorithm would turn to phase $k+1$ or terminate because of meeting the condition in the Line 2. That means once $N_{a}^{\text{ee},B}\geq\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$ holds, Algorithm 2 will never pull arm $a$ in the remaining phase $k$. We can assert $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\leq\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$. ∎
Lemma 30.
For $B$ such that $|B|=K$ and $\max_{a^{\prime}\in B}\mu_{a^{\prime}}<\mu_{0}$, Algorithm 2 must terminate no later than the end of phase $\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}$.
Proof of Lemma 30.
From the calculation in the proof of Lemma 29, we know
$t\geq\frac{112\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
$\Rightarrow$ $\hat{\mu}_{a,t}+\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}>\mu_{0}$
for arm $a\in B=[K]$. Notice that for any $a\in[K]$, we have
$k\geq\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil$
$\stackrel{{\scriptstyle\mu_{0}>\mu_{1}\geq\mu_{a}}}{{\Rightarrow}}$ $\beta_{k}\geq\frac{226}{\Delta_{a,0}^{2}},\log\frac{1}{\delta_{k}}\geq\frac{96}{\Delta_{a,0}^{2}}\log 3$
$\Rightarrow$ $\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{226\log\frac{2K}{\delta_{k}}}{\Delta_{a,0}^{2}},\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{226\log\log\frac{96}{\Delta_{a,0}^{2}}}{\Delta_{a,0}^{2}}$
$\Rightarrow$ $\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{113\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
$\stackrel{{\scriptstyle C>1}}{{\Rightarrow}}$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1\geq\frac{113\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}.$
That means at the end of phase $\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}$, we must have
$\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\log_{3}\frac{3}{\delta}\}}}{K})\leq\mu_{0}$
holds for all $a\in[K]$. Also, not hard to see $\delta_{\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}\leq\delta_{\lceil\log_{3}\frac{3}{\delta}\rceil}<\frac{\delta}{3}$. We can conclude Algorithm 2 must terminate by the condition in Line 2, 2. ∎
C.3 Proof of Theorem 2 and 3
The full proof requires multiple lemmas to support, which motivates us to use figure 1 to summarize the overall ideas. An arrow from A to B means Lemma A is used in the proof of Lemma/Theorem B. All the Lemmas inside the orange box are the properties of Algorithm 2, which are mainly adapted from Li and Cheung (2025).
Figure 1: Overall Idea of Full Proof of Theorem 2 and 3
Before illustrating the proof of Theorem 2 and 3, we first introduce a Lemma for the random permutation, Line 1 of Algorithm 1.
Lemma 31.
Given a qualified arm index $j\in[m]$, denote bracket index $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$. We have
$\Pr(b_{j}\geq\tilde{b})\leq\frac{1}{\exp\left(\frac{j|B_{\tilde{b}-1}|}{K}\right)},\forall\tilde{b}\in[\lceil\log_{2}K\rceil+1]$
Here we denote $B_{0}=\emptyset$ and $|B_{0}|=0$. Since we take $|B_{\tilde{b}}|=\min\{2^{\tilde{b}-1},K\}$ for $\tilde{b}\in[\lceil\log_{2}K\rceil+1]$, we can rewrite the conclusion as
$\Pr(b_{j}\geq\tilde{b})\leq\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{\tilde{b}-2}\rfloor}{K}\right)},\forall\tilde{b}\in[\lceil\log_{2}K\rceil+1].$
Proof.
If $\tilde{b}=1$, the conclusion is $\Pr(b_{j}\geq 1)\leq 1$, which is obviously true. We focus on the case that $2\leq\tilde{b}\leq\lceil\log_{2}K\rceil+1$. We have
[…]
$=$ $\Pr\Big(\{\sigma(i)\}_{i=1}^{2^{\tilde{b}-2}}\cap[j]=\emptyset\Big).$
The second line is from the fact that $\tilde{b}\in[\lceil\log_{2}K\rceil+1]$, which concludes $2^{\tilde{b}-2}\leq 2^{\lceil\log_{2}K\rceil-1}\leq 2^{\log_{2}K}=K$. We can conclude
$\Pr\Big(\{\sigma(i)\}_{i=1}^{2^{\tilde{b}-2}}\cap[j]=\emptyset\Big)$
$=$ $\prod_{i=1}^{2^{\tilde{b}-2}}\frac{K-j-i+1}{K-i+1}$
$\leq$ $(\frac{K-j}{K})^{2^{\tilde{b}-2}}$
$=$ $\left(\left(1-\frac{j}{K}\right)^{\frac{K}{j}}\right)^{\frac{j\cdot 2^{\tilde{b}-2}}{K}}$
$\leq$ $\exp\left(-\frac{j\cdot 2^{\tilde{b}-2}}{K}\right).$
Given the above preparation, we are ready to prove the Theorem 2 and 3. We first prove Theorem 2.
Proof of Theorem 2.
We first show $\Pr_{\nu}(\tau<+\infty)=1$, by applying Algorithm 1. This conclusion holds no matter $\nu$ is positive or negative. Consider the bracket $\lceil\log_{2}K\rceil+1$, which is indeed $[K]$. From Lemma 30 and Lemma 27, we know
$\Pr_{\nu}(\tau=+\infty)$
$\leq$ $\Pr_{\nu}(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)$
$\leq$ $\Pr_{\nu}(\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1}=+\infty)+\Pr_{\nu}(\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}=+\infty)$
holds for any instance $\nu$. By the Lemma 17, we know $\Pr_{\nu}(\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1}=+\infty)=\Pr_{\nu}(\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}=+\infty)=0$. That means $\Pr_{\nu}(\tau=+\infty)=0$, further $\Pr_{\nu}(\tau<+\infty)=1$. This suggest that $\hat{a}$ is well defined with probability 1, and we can assert $\Pr_{\nu}(\hat{a}\in[K]\cup\{\textsf{None}\})=1$ if $\nu$ is positive or negative.
Then, we suffice to show $\Pr_{\nu}(\mu_{\hat{a}}\leq\mu_{0}\text{ or }\mu_{\hat{a}}=\textsf{None})<\delta$ for a positive instance $\nu$ and $\Pr_{\nu}(\hat{a}\in[K])<\delta$ for a negative instance $\nu$. If $\nu$ is positive, we get
$\Pr_{\nu}(\mu_{\hat{a}}\leq\mu_{0}\text{ or }\hat{a}=\textsf{None})$
$\leq$ $\Pr_{\nu}(\mu_{\hat{a}}\leq\mu_{0})+\Pr_{\nu}(\hat{a}=\textsf{None})$
$\leq$ $\Pr_{\nu}(\text{alg copy }\lceil\log_{2}K\rceil+1\text{ outputs }\hat{a}=\textsf{None})+$
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\Pr_{\nu}(\text{alg copy }b\text{ outputs }\hat{a},s.t.\mu_{\hat{a}}\leq\mu_{0})$
$\leq$ $\frac{\pi^{2}}{6}\frac{\delta}{3}+\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{\pi^{2}}{6\cdot 5}\frac{\delta}{\lceil\log_{2}K\rceil+1}$
$\leq$ $\delta.$
The second last line is by Lemma 18 and 19.
If $\nu$ is negative, we get
$\Pr_{\nu}(\hat{a}\in[K])$
$\leq$ $\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\Pr_{\nu}(\text{alg copy }b\text{ outputs }\hat{a},s.t.\mu_{\hat{a}}\leq\mu_{0})$
$\leq$ $\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{\pi^{2}}{6\cdot 5}\frac{\delta}{\lceil\log_{2}K\rceil+1}$
$<$ $\delta.$
The second last line is by Lemma 18. ∎
To prove Theorem 3, we split the proof into two cases, $\nu$ is positive or negative.
Proof of Theorem 3, when $\nu$ is positive.
Fix arbitrary $j\in[m]$ and follow the notation in Lemma 31. Denote $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$. We have $\tau=\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\tau\mathds{1}(b_{j}=b)$. Then, we can turn to find the upper bound for each $\tau\mathds{1}(b_{j}=b)$.
From the algorithm design, we know
$\tau\mathds{1}(b_{j}=b)\leq$ $(\lceil\log_{2}K\rceil+1)\tau(B_{b})\mathds{1}(b_{j}=b)$
$\leq$ $(\lceil\log_{2}K\rceil+1)\Big(\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)+\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\Big).$
Then, we are going to derive an upper bound for $\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)$ and $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\Big)$ separately.
For $\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)$, we have
$\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)$
[…]
$\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\Bigg)$
[…]
$\underbrace{\sum_{a=1}^{K}\frac{29(C+1)^{2}\left(\log\frac{4K\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\mathds{1}(b_{j}=b,a\in B_{b})}_{\ddagger}.$
For $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)$, we have
$\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)$
[…]
$\leq$ $\underbrace{\frac{2(C+3)^{2}}{(C-1)^{2}}\mathds{1}(b_{j}=b)\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}|_{k=\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}_{\diamondsuit}.$
The last line is by the fact that $\frac{\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}}{\beta_{k-1}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k-1}}{\delta}}\geq 2$.
Then, to bound $\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)$ and $\mathbb{E}\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\Big)$, we suffice to consider the upper bound of the expectation of $\dagger,\ddagger,\diamondsuit$. For $\dagger$, we have
$\mathbb{E}\left[\mathds{1}(b_{j}=b)|B_{b}|\left(\beta_{\kappa^{\text{ee}}_{b}-1}\log\frac{4K}{\delta_{\kappa^{\text{ee}}_{b}-1}}+1\right)\right]$
$\leq$ $\Pr(b_{j}=b)|B_{b}|\mathbb{E}\left[\left(\beta_{\kappa^{\text{ee}}_{b}-1}\log\frac{4K}{\delta_{\kappa^{\text{ee}}_{b}-1}}+1\right)\right]$
$\leq$ $\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}\left(1+\sum_{k=1}^{+\infty}2^{k-1}\log(4K\cdot 3^{k-1})\frac{\pi^{2}}{6}\frac{1}{3^{k-1}}\right)$
$\leq$ $O(\log K)\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}.$
The second line is by the fact that $\kappa^{\text{ee}}_{b}$ is independent with $b_{j}$. The third line is by the Lemma 17, 31. For the term $\ddagger$, we have
$\mathbb{E}\left[\sum_{a=1}^{K}\frac{\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}+\log\log\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\mathds{1}(b_{j}=b)\mathds{1}(a\in B_{b})\right]$
$\leq$ $\sum_{a=1}^{K}\frac{\log\log\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\Pr(a\in B_{b},b_{j}=b)+$
$\sum_{a=1}^{K}\mathbb{E}\left[\frac{\log(4K)+(\kappa^{\text{ee}}_{b}+\kappa^{\text{et}}_{b}+\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil)\log 3}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right]\Pr(a\in B_{b},b_{j}=b)$
$\leq$ $\sum_{a=1}^{K}\frac{O\left((\log(4K)+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}})\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}+$
$\sum_{a=1}^{K}\frac{\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\cdot(\sum_{k=1}^{+\infty}k\frac{\pi^{2}}{6}\frac{1}{3^{k-1}}+\sum_{k=1}^{+\infty}k\frac{\pi^{2}}{6}\frac{\delta}{5^{k-1}}).$
The second step is by the fact that $\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\leq\frac{4}{\Delta_{j,0}^{2}}$ holds for all $a,j\in[K]$ and the third step is by the Lemma 17.
From Lemma 31, we know $\Pr(b_{j}=b)\leq\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}$. What’s more, we know $\Pr(a\in B_{b})=\frac{|B_{b}|}{K}=\frac{\min\{2^{b-1},K\}}{K}$ holds for all $a\in[K]$. Thus, we can conclude
$\mathbb{E}\left[\sum_{a=1}^{K}\frac{\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\log_{2}\frac{\Theta(1)}{\Delta_{j,0}^{2}}\}}}+\log\log\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\mathds{1}(b_{j}=b)\mathds{1}(a\in B_{b})\right]$
$\leq$ $\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\sum_{a=1}^{K}\frac{O\left((\log(4K)+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}})\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}+$
$\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\sum_{a=1}^{K}\frac{O(1)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$
$\leq$ $\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\sum_{a=1}^{K}\frac{O\left((\log(4K)+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}})\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}.$
For the expectation of $\diamondsuit$, we have
$\mathbb{E}\mathds{1}(b_{j}=b)\beta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}{\delta})$
$\leq$ $\Pr(b_{j}=b)O\left(\mathbb{E}\beta_{\kappa^{\text{ee}}_{b}}\log\frac{4K\alpha_{\kappa^{\text{ee}}_{b}}}{\delta}+\mathbb{E}\beta_{\kappa^{\text{et}}_{b}}\log\frac{4K\alpha_{\kappa^{\text{et}}_{b}}}{\delta}+\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$
$\leq$ $\Pr(b_{j}=b)O\left(\sum_{k=1}^{+\infty}\frac{\pi^{2}}{6}\frac{2^{k}}{3^{k}}\log\frac{4K\cdot 5^{k}}{\delta}+\sum_{k=1}^{+\infty}\frac{\pi^{2}}{6}\frac{2^{k}\delta}{5^{k}}\log\frac{4K\cdot 5^{k}}{\delta}+\frac{\log\frac{4K}{\delta}+\log_{2}\frac{\Theta(1)}{\Delta_{j,0}^{2}}\log 5}{\Delta_{j,0}^{2}}\right)$
$\leq$ $\Pr(b_{j}=b)O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$
Sum up all the upper bounds, we have
$\mathbb{E}\tau$
$\leq$ $O(\log K)\sum_{b=1}^{\lceil\log_{2}K\rceil+1}O(\log K)\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}+$
$O(\log K)\sum_{a=1}^{K}\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{O\left((\log(4K)+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}})\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}+$
$O(\log K)\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{(C+3)^{2}}{(C-1)^{2}}\Pr(b_{j}=b)O\left(\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$
By the Lemma 33, 34, we can conclude
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}\leq\frac{K}{j}\cdot O\Big((\log_{2}\frac{K}{j})+1\Big)$
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\leq O\Big(\frac{\log j}{j}\Big).$
Together with the fact $\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\Pr(b_{j}=b)=1$ and plug in default value $C=1.01$ specified by Algorithm 1, we have
$\mathbb{E}\tau$
$\leq$ $O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{K\log K\Big((\log_{2}\frac{K}{j})+1\Big)}{j}+\frac{\log j}{j}\sum_{a=1}^{K}\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)$
$\leq$ $O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log j\log K\Big((\log_{2}\frac{K}{j})+1\Big)}{j}\sum_{a=1}^{K}\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)$
$\leq$ $O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right).$
The second last line is by the assumption $\Delta_{i,j}<1$ for all $i,j\in[K]\cup\{0\}$. Notice the above upper bound holds for all $j\in[m]$, we can conclude
$\mathbb{E}\tau\leq O(\log K)\cdot O\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right).$
Proof of Theorem 3, when $\nu$ is negative.
Consider the bracket with size $K$, denoting this bracket as $B$. Similar to the proof of positive case, we split $\tau(B)=\tau^{\text{ee}}(B)+\tau^{\text{et}}(B)$ and prove an upper bound for $\mathbb{E}\tau^{\text{ee}}(B)$ and $\mathbb{E}\tau^{\text{et}}(B)$ separately. From Lemma 14, 29 and Lemma 30, we know
$\tau^{\text{ee}}(B)\leq$ $|Q|+\sum_{a=1}^{K}N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big)+\frac{O\left(\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}}+\log\log\frac{1}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
$=$ $K+\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)+\sum_{a=1}^{K}\frac{O\left(\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}}+\log\log\frac{1}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}.$
Taking expectation on both sides, we have
$\mathbb{E}\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)$
$\leq$ $\sum_{k=1}^{+\infty}|K|\beta_{k-1}\log\frac{4K}{\delta_{k-1}}\Pr(\kappa^{\text{ee}}_{B}=k)$
$\leq$ $\sum_{k=1}^{+\infty}|K|2^{k-1}\log(4K\cdot 3^{k-1})\frac{\pi^{2}}{6\cdot 3^{k-1}}$
$\leq$ $O(K\log K).$
The second last line is by the Lemma 17. In addition, we can also derive
$\mathbb{E}\log\frac{1}{\delta_{\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}}$
$\leq$ $\mathbb{E}\log\frac{1}{\delta_{\kappa^{\text{ee}}_{B}}}+\mathbb{E}\log\frac{1}{\delta_{\kappa^{\text{et}}_{B}}}+\mathbb{E}\log\frac{1}{\delta_{\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil}}+\mathbb{E}\log\frac{1}{\delta_{\lceil\log_{3}\frac{3}{\delta}\rceil}}$
$\leq$ $\sum_{k=1}^{+\infty}k\log 3\Pr(\kappa^{\text{ee}}_{B}=k)+\sum_{k=1}^{+\infty}k\log 3\Pr(\kappa^{\text{et}}_{B}=k)+O\left(\log_{2}\frac{1}{\Delta_{1,0}^{2}}\right)+\log\frac{9}{\delta}$
$\leq$ $\sum_{k=1}^{+\infty}\frac{\pi^{2}}{6}\frac{k}{3^{k-1}}\log 3\Pr(\kappa^{\text{ee}}_{B}=k)+\sum_{k=1}^{+\infty}\frac{\pi^{2}}{6}\frac{k\delta}{5^{k-1}}\log 3+O\left(\log_{2}\frac{1}{\Delta_{1,0}^{2}}\right)+\log\frac{3}{\delta}$
$\leq$ $O\left(\log_{2}\frac{1}{\Delta_{1,0}^{2}}+\log\frac{1}{\delta}\right).$
The second last line is still by the Lemma 17. The above two inequalities suggests that
$\mathbb{E}\tau^{\text{ee}}(B)\leq$ $O(K\log K)+\sum_{a=1}^{K}O\left(\frac{\log\frac{12K(\lceil\log_{2}K\rceil+1)}{\delta}+\log\frac{1}{\Delta_{0,1}^{2}}}{\Delta_{a,0}^{2}}\right)$
$\leq$ $\sum_{a=1}^{K}O\left(\frac{\log\frac{12K(\lceil\log_{2}K\rceil+1)}{\delta}+\log\frac{1}{\Delta_{0,1}^{2}}}{\Delta_{a,0}^{2}}\right).$
Meanwhile, for $\mathbb{E}\tau^{\text{et}}(B)$, we can utilize Lemma 28 to conclude
$\mathbb{E}\tau^{\text{et}}(B)\leq$ $\mathbb{E}\sum_{k=1}^{\kappa^{\text{ee}}_{B}-1}O(1)\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}$
$\leq$ $\mathbb{E}O(1)\beta_{\kappa^{\text{ee}}_{B}-1}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{\kappa^{\text{ee}}_{B}-1}}{\delta}$
$\leq$ $O(1)\sum_{k=1}^{+\infty}\beta_{k-1}\frac{\pi^{2}}{6}\delta_{k-1}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k-1}}{\delta}$
$\leq$ $O\left(\log\frac{4K}{\delta}\right).$
Combining the upper bound of $\mathbb{E}\tau^{\text{ee}}(B)$ and $\mathbb{E}\tau^{\text{et}}(B)$, we have
$\mathbb{E}\tau\leq O(\log K)\mathbb{E}\tau(B)\leq O(\log K)O\left(\sum_{a=1}^{K}\frac{\log\frac{K}{\delta}+\log\frac{1}{\Delta_{0,1}^{2}}}{\Delta_{a,0}^{2}}\right)$
Appendix D Technical Lemma
This section is for the proof of inequalities.
Lemma 32.
Denote $\{Z_{s}\}_{s=1}^{+\infty}$ is a group of random variable fulfilling that $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(0,\Delta^{2})$ for some $\Delta>0$. Then, we have
$\Pr(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}>z)\leq\exp\left(-\frac{z^{2}}{2n\Delta^{2}}\right).$
Proof of Lemma 32.
Notice that for $\lambda=\frac{z}{n\Delta^{2}}$, we have
$\Pr\big(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}\geq s\big)$
$=$ $\Pr\big(\max_{1\leq t\leq n}\lambda\big(\sum_{s=1}^{t}Z_{s}\big)\geq\lambda s\big)$
$=$ $\Pr\left(\max_{1\leq t\leq n}\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\geq\exp\left(\lambda s\right)\right).$
Easy to see $\{\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\}_{t=1}^{+\infty}$ is a submartingale, as
$\mathbb{E}[\exp\left(\lambda\big(Z_{t}\big)\right)|z_{1},\cdots,z_{t-1}]$
$\stackrel{{\scriptstyle\text{Jensen's Inequality}}}{{\geq}}$ $\exp\left(\lambda\mathbb{E}[\big(Z_{t}\big)|Z_{1},\cdots,Z_{t-1}]\right)$
$=$ $\exp(0)=1.$
Then, we can conclude
$\Pr(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}>z)$
$\leq$ $\frac{\mathbb{E}_{\nu^{\prime}}\prod_{t=1}^{n}\exp(\lambda(Z_{t}))}{\exp(\lambda s)}$
$=$ $\frac{\exp(\frac{n\lambda^{2}\Delta^{2}}{2})}{\exp(\lambda s)}$
$=$ $\exp(-\frac{s^{2}}{2n\Delta^{2}}).$
Lemma 33.
For any $K\in\mathbb{N}$, $K\geq 2$, $j\in[K]$ we have
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}\leq\frac{K}{j}\cdot O\Big((\log_{2}\frac{K}{j})+1\Big)$
Proof.
We can conduct direct calculation.
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}$
$\stackrel{{\scriptstyle K\geq 2}}{{=}}$ $1+\frac{2}{\exp(\frac{j}{K})}+\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}$
$\leq$ $1+\frac{2}{\exp(\frac{j}{K})}+\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{2^{b-1}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$=$ $1+\frac{2}{\exp(\frac{j}{K})}+\frac{2K}{j}\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $1+\frac{2K}{j}+\frac{2K}{j}\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}.$
Then, we suffice to prove $\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}\leq O\left((\log_{2}\frac{K}{j})+1\right)$.
From the fact that $\frac{d\frac{x}{\exp(x)}}{dx}=e^{-x}-xe^{-x}=(1-x)e^{-x}$, we know $\frac{x}{\exp(x)}$ monotonically increases in the interval $(0,1)$ and decreases in the interval $(1,+\infty)$. We have
$\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $\sum_{b=3}^{\lfloor\log_{2}\frac{K}{j}\rfloor+3}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}+\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $(\lfloor\log_{2}\frac{K}{j}\rfloor+1)\frac{x}{e^{x}}|_{x=1}+\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $(\lfloor\log_{2}\frac{K}{j}\rfloor+1)\frac{1}{e}+\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}.$
The second last line is by that function $\frac{x}{e^{x}}$ achieves its maximum at $x=1$. The remaining work is to prove $\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}\leq O(1)$.
Notice that for $b\geq\lceil\log_{2}\frac{K}{j}\rceil+3$, we have $\frac{j2^{b-2}}{K}\geq\frac{j2^{\log_{2}\frac{K}{j}+1}}{K}\geq 2$. Since we can conclude $\frac{x}{e^{x}}\geq 2\cdot\frac{2x}{e^{2x}}$ holds for all $x\geq 2$, we can conclude
$\frac{\frac{j\cdot 2^{b-1}}{K}}{\exp\left(\frac{j\cdot 2^{b-1}}{K}\right)}/\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}\leq\frac{1}{2}$
holds for all $b\geq\lceil\log_{2}\frac{K}{j}\rceil+3$. Thus, we have
$\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{+\infty}2^{-(b-\lceil\log_{2}\frac{K}{j}\rceil-3)}\frac{\frac{j\cdot 2^{\lceil\log_{2}\frac{K}{j}\rceil+3-2}}{K}}{\exp\left(\frac{j\cdot 2^{\lceil\log_{2}\frac{K}{j}\rceil+3-2}}{K}\right)}$
$\leq$ $\frac{1}{e}\sum_{b=0}^{+\infty}2^{-b}$
$\leq$ $O(1).$
That means we complete the proof. ∎
Lemma 34.
For any $K\in\mathbb{N}$, $K\geq 2$, $j\in[K]$ we have
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\leq O\Big(\frac{\log j}{j}\Big).$
Proof.
We can conduct direct calculation.
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}$
$\leq$ $\frac{1}{K}+\sum_{b=2}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)},\frac{2}{j}\frac{j\cdot 2^{b-2}}{K}\}$
$\leq$ $\frac{1}{K}+\sum_{b=2}^{\lceil\log_{2}\frac{K\log j}{j}\rceil+2}\frac{2}{j}\frac{j\cdot 2^{b-2}}{K}+\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}.$
We prove an upper bound for $\sum_{b=2}^{\lceil\log_{2}\frac{K\log j}{j}\rceil+2}\frac{2}{j}\frac{j\cdot 2^{b-2}}{K}$ and $\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$ separately.
First, we have
$\sum_{b=2}^{\lceil\log_{2}\frac{K\log j}{j}\rceil+2}\frac{2}{j}\frac{j\cdot 2^{b-2}}{K}$
$\leq$ $\frac{2}{j}\cdot 2\cdot\frac{j2^{b-2}}{K}|_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+2}$
$\leq$ $\frac{2}{j}\cdot 2\cdot\frac{j2^{1+\log_{2}\frac{K\log j}{j}}}{K}$
$=$ $\frac{8\log j}{j}.$
For $\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$, we first notice that when $b\geq\lceil\log_{2}\frac{K\log j}{j}\rceil+3$, we have
$\frac{j2^{b-2}}{K}\geq\frac{j2^{\lceil\log_{2}\frac{K\log j}{j}\rceil+3-2}}{K}\geq\frac{j2^{\log_{2}\frac{K\log j}{j}+1}}{K}\geq 2.$
Further, we can conclude for $b\geq\lceil\log_{2}\frac{K\log j}{j}\rceil+3$, we have $\frac{j2^{b-2}}{K}+1\leq\frac{j2^{b-2}}{K}\cdot 2=\frac{j2^{b-1}}{K}$. Thus, we have
$\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{+\infty}\frac{1}{\exp(b-\lceil\log_{2}\frac{K\log j}{j}\rceil-3)}\frac{1}{\exp\left(\frac{j\cdot 2^{\lceil\log_{2}\frac{K\log j}{j}\rceil+3-2}}{K}\right)}$
$\leq$ $\frac{2}{\exp\left(\frac{j\cdot 2^{\lceil\log_{2}\frac{K\log j}{j}\rceil+3-2}}{K}\right)}$
$\leq$ $\frac{2}{\exp\left(\frac{j\cdot 2^{\log_{2}\frac{K\log j}{j}+1}}{K}\right)}$
$=$ $\frac{2}{\exp\left(\frac{j\cdot 2\cdot\frac{K\log j}{j}}{K}\right)}$
$=$ $\frac{2}{\exp\left(2\log j\right)}$
$<$ $\frac{2}{j}.$
In summary, we have proved
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}$
$\leq$ $\frac{1}{K}+\frac{8\log j}{j}+\frac{2}{j}$
$\leq$ $O\Big(\frac{\log j}{j}\Big).$
Lemma 35 (Adapted Lemma 3 in Jamieson et al. (2014), Lemma D.3 in Li and Cheung (2025)).
Denote $\{X_{i}\}_{i=1}^{+\infty}$ as i.i.d $\sigma^{2}$-subgaussian random variable with true mean reward $\mu=0$. For any $\delta\in(0,1)$, we have
$\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2(\log_{2}2^{\lceil\log_{2}t\rceil^{+}})^{2}}{\delta}}\right)<\frac{\pi^{2}}{6}\delta.$
Or equivalently,
$\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}\right)<\frac{\pi^{2}}{6}\delta.$
Proof of Lemma 35.
Define $u_{k}=2^{k}$, $k\geq 1$. Define $x=\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}$, $S_{t}=\sum_{i=1}^{t}X_{i}$ and the event
$E_{k}=\left\{\max_{1\leq t\leq u_{k}}S_{t}>\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}\right\}\cup\left\{\min_{1\leq t\leq u_{k}}S_{t}<-\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}\right\},$
For $\lambda>0$, notice that
$\mathbb{E}\left[\exp(\lambda(\sum_{s=1}^{t}X_{s}))|X_{1},\cdots,X_{t-1}\right]$
$=$ $\exp(\lambda(\sum_{s=1}^{t-1}X_{s}))\mathbb{E}\exp(\lambda X_{t})$
$\geq$ $\exp(\lambda(\sum_{s=1}^{t-1}X_{s}))\exp(\mathbb{E}\lambda X_{t})$
$=$ $\exp(\lambda(\sum_{s=1}^{t-1}X_{s})).$
Take $\lambda=\frac{x}{u_{k}\sigma^{2}}$, we can conclude $\{\exp(\lambda S_{t})\}$ is a submartingale. Then,
$\Pr\left(\max_{1\leq t\leq u_{k}}S_{t}\geq x\right)$
$=$ $\Pr\left(\max_{1\leq t\leq u_{k}}\exp(\lambda S_{t})>\exp\left(\lambda x\right)\right)$
$\stackrel{{\scriptstyle*}}{{\leq}}$ $\frac{\mathbb{E}\exp(\lambda S_{u_{k}})}{\exp\left(\lambda x\right)}$
$\leq$ $\frac{\exp(\frac{u_{k}\lambda^{2}\sigma^{2}}{2})}{\exp\left(\lambda x\right)}$
$\stackrel{{\scriptstyle\lambda=\frac{x}{u_{k}\sigma^{2}}}}{{=}}$ $\exp(-\frac{x^{2}}{2u_{k}\sigma^{2}}).$
Step * is by the maximal inequality for the submartingale. Take $x=\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}$, we have
$\Pr\left(\max_{1\leq t\leq u_{k}}S_{t}\geq\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}\right)\leq\exp\left(-\log\frac{2(\log_{2}u_{k})^{2}}{\delta}\right)=\frac{\delta}{2(\log_{2}u_{k})^{2}}=\frac{\delta}{2k^{2}}$
For the part of $\Pr\left(\min_{1\leq t\leq u_{k}}S_{t}<-x\right)$, the proof is similar. We can conclude $\Pr(E_{k})\leq\frac{\delta}{k^{2}}$ and further $\Pr(\cup_{k=1}^{+\infty}E_{k})\leq\frac{\pi^{2}\delta}{6}$.
Thus,
$\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}\max\{2^{\lceil\log_{2}t\rceil},2\}\log\frac{2(\log_{2}\max\{2^{\lceil\log_{2}t\rceil},2\})^{2}}{\delta}}\right)$
$\leq$ $\Pr\left(\exists k,\max_{1\leq t^{\prime}\leq u_{k}}|\sum_{s=1}^{t^{\prime}}X_{s}|\geq\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}\right)$
$\leq$ $\Pr(\cup_{k=1}^{+\infty}E_{k})\leq\frac{\pi^{2}\delta}{6}.$
Some Comments are as follows.
  • We can similarly prove $\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2\pi^{2}(\log_{2}2^{\lceil\log_{2}t\rceil^{+}})^{2}}{6\delta}}\right)<\delta$ holds for all $\delta\in(0,1)$.
  • Since $\lceil\log_{2}t\rceil^{+}\leq 1+\log_{2}t$, we have
    $\frac{\pi^{2}\delta}{6}\geq$ $\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}2^{\max\{\lceil\log_{2}t\rceil,1\}}\log\frac{2(\log_{2}2^{\lceil\log_{2}t\rceil})^{2}}{\delta}}\right)$
    $\geq$ $\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{4\sigma^{2}t\log\frac{2(\log_{2}2t)^{2}}{\delta}}\right)$
Lemma 36 (Lemma D.2 in Li and Cheung (2025)).
For any $\Delta\in(0,1],K\geq 2,\delta\in(0,\frac{1}{2}],C\geq 1$, we can conclude
$t>\frac{28C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{16C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)\right)}{\Delta^{2}}$
$\Rightarrow$ $C\sqrt{\frac{4\log\frac{2K(\log_{2}2t)^{2}}{\delta}}{t}}<\Delta$
Proof.
By simple calculation, we can derive
$t>\frac{28C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{16C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)\right)}{\Delta^{2}}$
$\Leftrightarrow$ $2t>\frac{56C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)\right)}{\Delta^{2}}$
$\Leftrightarrow$ $2t>\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)\right)+32C^{2}\log(\frac{2K}{\delta})}{\Delta^{2}}$
$\stackrel{{\scriptstyle\log(x+y)\leq\log x+\log y,\forall x,y\geq 2}}{{\Rightarrow}}$ $2t>\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)+\frac{2K}{\delta}\right)}{\Delta^{2}}$
$\Rightarrow$ $2t>\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)+\log\left(\log\frac{2K}{\delta}\right)\right)}{\Delta^{2}}$
$\Leftrightarrow$ $2t>\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\log\left(\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}\right)}{\Delta^{2}}$
[…]
$\Rightarrow$ $2t>\frac{8C^{2}\log\frac{2K}{\delta}+16C^{2}\log\log_{2}e+16C^{2}\log\log(2t)}{\Delta^{2}}$
$\Leftrightarrow$ $t>\frac{4C^{2}\log\frac{2K}{\delta}+8C^{2}\log\frac{\log(2t)}{\log 2}}{\Delta^{2}}$
$\Leftrightarrow$ $t>\frac{4C^{2}\log\frac{2K}{\delta}+8C^{2}\log(\log_{2}2t)}{\Delta^{2}}$
$\Leftrightarrow$ $C\sqrt{\frac{4\log\frac{2K(\log_{2}2t)^{2}}{\delta}}{t}}<\Delta.$
Lemma 37.
For any $b\geq a\geq 0$,
  • If $b\geq e^{2}$, we have $x\geq b+2a\log\log b\Rightarrow x\geq a\log\log(x)+b$.
  • If $b,a\geq e$, we have $e\leq x\leq b+a\log\log b\Rightarrow x<a\log\log(x)+b$.
The second inequality also implies $x\geq a\log\log(x)+b\Rightarrow x\geq b+a\log\log b$.
Proof.
We prove the first claim. Easy to see $\frac{d(x-a\log\log x-b)}{dx}=1-\frac{a}{x\log x}=\frac{x\log x-a}{x\log x}$. Take $x_{0}=b+2a\log\log b$, then $x_{0}\log x_{0}>x_{0}\log(e^{2}+2a\log\log e^{2})>x_{0}\log e^{2}=2x_{0}>a$. Thus $x-a\log\log(x)-b$ increases in the interval $(x_{0},+\infty)$. On the other hand, easy to check
$x_{0}-a\log\log x_{0}-b$
$=$ $b+2a\log\log b-a\log\log(b+2a\log\log b)-b$
$=$ $2a\log\log b-a\log\log(b+2a\log\log b)$
$=$ $a\left(2\log\log b-\log\log(b+2a\log\log b)\right)$
$=$ $a\left(\log(\log b)^{2}-\log\log(b+2a\log\log b)\right)$
$=$ $a\log\frac{(\log b)^{2}}{\log(b+2a\log\log b)}$
$\geq$ $a\log\frac{(\log b)^{2}}{\log(b+2b\log\log b)}$
$=$ $a\log\frac{(\log b)^{2}}{\log b+\log(1+2\log\log b)},$
Notice that
$(\log b)^{2}-\log b-\log(1+2\log\log b)$
$=$ $\log b(\log b-1)-\log(1+2\log\log b)$
$\stackrel{{\scriptstyle b\geq e^{2}}}{{\geq}}$ $\log b-\log(1+2\log\log b)$
$\geq$ $\log b-2\log\log b$
$\stackrel{{\scriptstyle x>2\log x,\forall x>0}}{{>}}$ $0,$
we can conclude $x_{0}-a\log\log x_{0}-b>0$ holds for all $x\geq b+2a\log\log b$.
Then we turn to prove the second claim. As $\frac{d(x-a\log\log x-b)}{dx}=1-\frac{a}{x\log x}=\frac{x\log x-a}{x\log x}$, we know there is at most 1 zero point of $\frac{x\log x-a}{x\log x}$ in the interval $(e,b+a\log\log b)$. Thus,
$\max_{e\leq x\leq b+a\log\log b}x-a\log\log x-b=\max\{x-a\log\log x-b|_{x=e},x-a\log\log x-b|_{x=b+a\log\log b}\}.$
Easy to see
$e-a\log\log e-b=e-b<0,$
and
$b+a\log\log b-a\log\log(b+a\log\log b)-b$
$=$ $a\log\log b-a\log\log(b+a\log\log b)$
$<$ $a\log\log b-a\log\log(b)$
$=$ $0.$
That means $\max\limits_{e\leq x\leq b+a\log\log b}x-a\log\log x-b<0$. The second conclusion is done. ∎